### v1

# Classifying representations of finite classical groups of Lie type of dimension up to $\ell^{4}$

Abstract.
In this article we classify the irreducible representations of finite classical groups $G$ of Lie type of rank $\ell$ with dimension at most a constant proportional to $\ell^{4}$ according to whether $G$ is of type $A_{\ell}$ or not. The dimensions of those representations are explicitly computed by hand. We conclude the work by addressing certain cases of the inverse Galois problem via the foregoing classification.
2020 Mathematics Subject Classification. 20C33, 20G40, 11F80.
1. Introduction
It is well-known the multiples bridges among representation theory with other theories in mathematics. We are specially concerned about the irreducible representations of finite classical groups of Lie type and its connections with the large image conjecture for automorphic Galois representations and the inverse Galois problem. More precisely, let $G_{\mathbb{Q}}:=\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ be the absolute Galois group of $\mathbb{Q}$ and $\mathbb{A}_{\mathbb{Q}}$ the ring of adeles of $\mathbb{Q}$ and $\pi$ be a cuspidal algebraic regular automorphic representations of $\operatorname{GL}_{n}(\mathbb{A}_{\mathbb{Q}})$. According to global Langlands correspondence [5], for each prime $p$ there should exist a $p$-adic Galois representation $\rho_{\pi,p}:G_{\mathbb{Q}}\rightarrow\operatorname{GL}_{n}(\overline{\mathbb{Q}}_{p})$ associated to $\pi$. A folklore conjecture (usually known as the large image conjecture) predicts that the image of the proyectivization $\overline{\rho}^{proj}_{\pi,p}:G_{\mathbb{Q}}\rightarrow\operatorname{PGL}_{n}(\overline{\mathbb{F}}_{p})$ of the residual representation $\overline{\rho}_{\pi,p}:G_{\mathbb{Q}}\rightarrow\operatorname{GL}_{n}(\overline{\mathbb{F}}_{p})$ of $\rho_{\pi,p}$ should be as large as possible for almost all primes $p$ (i.e., all but finitely many), except when an automorphic reason impede it. For instance, when $\pi$ is polarizable [2, §2.1], this means that the image of $\overline{\rho}^{proj}_{\pi,p}$ should contain either $\operatorname{PSp}_{n}(\mathbb{F}_{p^{s}})$ or $\operatorname{P\Omega}^{\pm}_{n}(\mathbb{F}_{p^{s}})$ (for some positive integer $s$) for almost all $p$. This conjecture is, in general, known in very few cases ($n\leq 6$) and the proofs in all known cases when $\pi$ is polarizable are done by considering all possible images of $\overline{\rho}^{proj}_{\pi,p}$ given by the classification of the maximal subgroups of $\operatorname{PSp}_{n}(\mathbb{F}_{p^{s}})$ or $\operatorname{P\Omega}^{\pm}_{n}(\mathbb{F}_{p^{s}})$see [15] for more details. As we will see at the end of this paper, this conjecture has important consequences on the Inverse Galois problem for finite groups of Lie type.
[…]
We would like to point out that we are assuming $G$ of simply connected type, which ensures a correspondence between representations of the group $G$ and representations of its Lie algebra $\mathfrak{g}$ for almost all $p$. Therefore, through this paper, we will consider $p$ sufficiently large in order to account with such a correspondence. The main result of this paper splits into three theorems which tackle the classification of irreducible representations of $G$ of dimension less or equal to $\ell^{4}/16$ or $\ell^{4}$ according to whether the group is of type $A_{\ell}$ or not. Thanks to the work of Chevalley, we know that each irreducible $G(\overline{\mathbb{F}}_{p})$-module $V$ is determined (up to isomorphism) by its highest weight and for each dominant weight $\lambda$ there is an irreducible $G(\overline{\mathbb{F}}_{p})$-module $V(\lambda)$ with highest weight $\lambda$. Thereby, classifying those irreducible representations is equivalent to determine all dominant weights $\lambda$ such that the dimension of the irreducible $G(\overline{\mathbb{F}}_{p})$-module $V(\lambda)$ of highest weight $\lambda$ satisfies the foregoing condition-we will call them admissible weights. Determining this class of weights will imply to compute diverse dimensions and consequently to handle a considerable number of multiplicities. We then will present certain examples with these calculations first to illustrate admissible and not admissible weights. Nevertheless, one of the key motivations to give these examples is to reminisce about the computing of multiplicities and its recursive nature since we will have to compute longer multiplicities and dimensions along the work. In addition, a second good motivation is that those examples are placed in such a manner to contribute in the proofs of our main results, helping to be more readable such demonstrations.
The classification of admissible weights splits into three stages according to $G$ is of type $A_{\ell}$, $B_{\ell}$ or of type $C_{\ell}$ and $D_{\ell}$. Our strategy to this end will be to analyze the behavior of diverse multiplicities of weights $\mu\preceq\lambda$ to study the dimension of $V(\lambda)$ whereby we obtain a good list of admissible weights in order to show that this list exhaust the admissible ones. We therefore commence by applying successfully the mentioned strategy to $G$ of type $A_{\ell}$ as starting case. Later on, we tackle the case when $G$ is of type $B_{\ell}$, where we may point out that determining the admissible weights certainly presents additional challenges. Indeed, computing the required multiplicities to study admissibility of weights implies much more efforts than the case $A_{\ell}$. In particular, these calculations turn out to be very laborious with extensive computations to use the Freudenthal formula to determine multiplicities at hand. We are able to carry out all those computations and find a good list of admissible weights. Once we have ended that, obtaining good lists for types $C_{\ell}$, $D_{\ell}$ will follow more natural.
The second stage of the strategy also will be a subtle task in the sense that from an arbitrary admissible weight we will have to prove this one can be reduced to a one of our good list, i.e., the list exhausts the admissible ones which gives rise to the desired classification. We here continue case-by-case tackling first the type $A_{\ell}$ and subsequently the case of type $B_{\ell}$. This work takes a significant time and whence the cases $C_{\ell}$ and $D_{\ell}$ will be deduced directly. Hence we only outline the proof by refining the arguments according to the case and by proving the admissibility of the new weight that appears when $G$ is of type $C_{\ell}$ and $D_{\ell}$. At this point, we would like to observe that the required multiplicities in these cases are listed in two tables of the Appendix, where we bring all information to compute them, thus avoiding repetitive arguments and calculations.
We conclude the Introduction by giving the structure of the paper.
This article begins with a brief background on representation theory of finite groups of Lie type in §2. Here we introduce the root systems and fix the orders that we will be used through the work. Also we recall the Freudenthal multiplicity formula and the expression to compute dimension of the $G(\overline{\mathbb{F}}_{p})$-module $V(\lambda)$ with highest weight $\lambda$. In §3 we introduce the concept of admissible weight $\lambda$ when $G$ is of type $A_{\ell}$ and obtain a good list -Proposition 3.6- which is the first part toward the classification of admissible weights of $G$. Later on, §4 is dedicated completely to Theorem 4.1, where the classification of the admissible weights is obtained when $G$ is of type $A_{\ell}$. In §5 we tackle our aim for groups $G$ of type $B_{\ell}$, $C_{\ell}$ and $D_{\ell}$. In the first case we prove a series of lemmas on required multiplicities to obtain a list of admissible weights -Proposition 5.6. Thus, we may show such a classification for $G$ of type $B_{\ell}$ in Theorem 5.9. While Theorem 5.10 determines all the admissible weights for the remainder two types by appealing to Appendix -Tables 1 and 2- for the required multiplicities. We present the last two results of the article in §6 by providing applications of our classification to large image conjecture for automorphic Galois representation in Theorem 6.2 and on the inverse Galois problem in Corollary 6.3. We conclude our work with an appendix §7 where we give the tables with the all required multiplicities when $G$ is of type $C_{\ell}$ or $D_{\ell}$.
2. Preliminaries on representations of finite simple groups of Lie type
In this section we will recall some results about algebraic groups and their defining characteristic representations that we will need in this paper.
2.1. Root systems
Let $G$ be a connected reductive simple algebraic group over $\overline{\mathbb{F}}_{p}$ of simply connected type and rank $\ell$. Let $T$ be a maximal torus of $G$, $X\cong\mathbb{Z}^{\ell}$ its character group and $Y\cong\mathbb{Z}^{\ell}$ its co-character group. Let $\{\alpha_{1},\ldots,\alpha_{\ell}\}\subset X$ be a set of simple roots for $G$ with respect to $T$ and $\Delta$ its associated root system. Let $\alpha_{i}^{\vee}\in Y$ the coroot corresponding to $\alpha_{i}$ for each $1\leq i\leq\ell$. Viewing $X\otimes\mathbb{R}$ as Euclidean space we can define the fundamental weights $\lambda_{1}\ldots,\lambda_{\ell}\in X\otimes\mathbb{R}$ as the dual basis of $\alpha_{1}^{\vee},\ldots,\alpha_{\ell}^{\vee}$.
For our purposes in accordance with [11] we shall use the opposite order from the one given in [3, 8]. Thus, the Dynkin diagrams for classical groups are given as follows:
We also recall that each $\alpha\in\Delta$ determines a reflection $\sigma_{\alpha}$ by means of
$\sigma_{\alpha}(\beta)=(\alpha,\beta),\quad\quad\beta\in E,$
where $(E,(-,-))$ is the underlying inner product real vector space of finite dimensional associated to $\Delta$. The Weyl group $\mathcal{W}$ of $\Delta$ is the subgroup of the orthogonal group on $E$ generated by the reflections $\sigma_{\alpha_{i}}$. Thus one observes that this group acts on $\Delta$ by permutation. In fact, $\mathcal{W}$ is generated by the reflections $\sigma_{\alpha_{i}}$ with $1\leq i\leq\ell$ by appealing to [9, Prop. 2.62].
Furthermore, one obtains a partial ordering in $X$ by declaring that a weight $\lambda$ is smaller than a weight $\lambda^{\prime}$ (denoted by $\lambda\prec\lambda^{\prime}$) if and only if $\lambda^{\prime}-\lambda$ is a non-negative linear combination of simple roots. We shall also use $\preceq$ in an obvious way. In addition, a weight $\lambda\in X$ is dominant if it is a non-negative linear combination of the fundamental weights. The group $\mathcal{W}$ acts on the set of the dominant weights by means of its action on the simple roots $\alpha_{1},\cdots,\alpha_{\ell}$. Given a dominant weight $\lambda$, we will denote by $\operatorname{Orb}_{\mathcal{W}}(\lambda)$ and $\rm{Stab}_{\mathcal{W}}(\lambda)$ the orbit and the stabilizer of $\lambda$ under the $\mathcal{W}$-action, respectively.
It is worth recalling that the stabilizer subgroup $\rm{Stab}_{\mathcal{W}}(\lambda)$ of a dominant weight $\lambda=a_{1}\lambda_{1}+a_{2}\lambda_{2}+\cdots+a_{\ell}\lambda_{\ell}$ is the parabolic subgroup of $G$ generated by the reflections along the simple roots $\alpha_{i}$ such that $a_{i}=0$.
2.2. Representations
Let $V$ be a finite-dimensional $G(\overline{\mathbb{F}}_{p})$-module. Considering this as $T$-module, we have a decomposition $V=\bigoplus_{\lambda\in X}V_{\lambda}$ into weight spaces $V_{\lambda}=\{v\in V:vt=\lambda(t)v\mbox{ for all }t\in T\}$. Then, the set of weights $\lambda\in X$ with $V_{\lambda}\neq\{0\}$ is called the set of weights of $V$. This set is a union of $\mathcal{W}$-orbits.
As we pointed out in the Introduction, from a result of Chevalley (see Section 31.3 of [8]) we have that, if $V$ is irreducible, then the set of weights of $V$ contains a unique element $\lambda$ such that for all weight $\lambda^{\prime}$ of $V$ we have $\lambda^{\prime}\prec\lambda$. This $\lambda$ is called the highest weight of $V$. This is dominant and $\dim(V_{\lambda})=1$. Moreover, we have that each irreducible $G(\overline{\mathbb{F}}_{p})$-module $V$ is determined (up to isomorphism) by its highest weight and for each dominant weight $\lambda\in X$ there is an irreducible $G(\overline{\mathbb{F}}_{p})$-module $V(\lambda)$ with highest weight $\lambda$.
For example, when $G$ is of $A_{\ell}$ type, the standard representation $V=V(\lambda)$ has highest weight $\lambda=\lambda_{\ell}$ and dimension $\ell+1$. Moreover, the anti-symmetric tensor product $\bigwedge^{k}(V)=V(\lambda)$ is irreducible with highest weight $\lambda=\lambda_{\ell-k+1}+\cdots+\lambda_{\ell}$ and dimension $\binom{\ell+1}{k}$. More generally, the irreducible representation $V(\lambda)$, with highest weight $\lambda=a_{1}\lambda_{1}+(a_{1}+a_{2})\lambda_{2}+\cdots+(a_{1}+\cdots+a_{\ell})\lambda_{\ell}$, is contained in
$\operatorname{Sym}^{a_{1}}(\Lambda^{\ell}V)\otimes\cdots\otimes\operatorname{Sym}^{a_{\ell-1}}(\Lambda^{2}V)\otimes\operatorname{Sym}^{a_{\ell}}(V)$
by Weyl’s Construction [7, Lecture 15]. In particular, taking $a_{\ell}\neq 0$ and $a_{i}=0$, $i=1,\cdots,\ell-1$, we have that $V(a_{\ell}\lambda_{\ell})$ is contained in the $a_{\ell}$-symmetric power $\operatorname{Sym}^{a_{\ell}}(V)$ of $V$, which has dimension $\binom{\ell+a_{\ell}}{a_{\ell}}$. In fact, we will prove that $\operatorname{Sym}^{a_{\ell}}(V)$ is irreducible by computing the dimension of $V(a_{\ell}\lambda_{\ell})$. In general, when $G$ is of type $B_{\ell}$, $C_{\ell}$ and $D_{\ell}$, we can only guarantee that $V(k\lambda_{\ell})\subseteq\operatorname{Sym}^{k}(V)$ and $V(\lambda_{\ell-k+1}+\cdots+\lambda_{\ell})\subseteq\Lambda^{k}(V)$. Through this paper, we also compute the exact dimensions of these irreducible representations $V(\lambda)$ for $k=2,3,4$ by computing multiplicities as follows.
2.3. Multiplicity and dimension
First, we introduce a special weight which appears in computing of multiplicities as we will see here below. Specifically, the strongly dominant weight $\delta$ is defined by
$\delta:=\frac{1}{2}\sum_{\alpha\succ 0}\alpha=\sum_{i=1}^{\ell}\lambda_{i},$
following [8, §10.2 and §13.3].
Now, let $V(\lambda)$ be a finite-dimensional $G(\overline{\mathbb{F}}_{p})$-module with highest weight $\lambda$ and $\mu\preceq\lambda$ be a dominant weight. We define the multiplicity of $\mu$ in $V(\lambda)$ as
$m_{\lambda}(\mu)=\dim V(\lambda)_{\mu}$
which can be calculated by using the recursive Freudenthal multiplicity formula [8, §22.3]:
$m_{\lambda}(\mu)=\frac{2\sum_{\alpha\in\Phi_{+}}\sum_{i=1}^{\infty}(\mu+i\alpha,\alpha)m_{\lambda}(\mu+i\alpha)}{\parallel\lambda+\delta\parallel^{2}+\parallel\mu+\delta\parallel^{2}}=\frac{2\sum_{\alpha\in\Phi_{+}}\sum_{i=1}^{\infty}(\mu+i\alpha,\alpha)m_{\lambda}(\mu+i\alpha)}{(\lambda,\lambda)-(\mu,\mu)+2(\lambda-\mu,\delta)}$
where $\Phi_{+}$ denotes the set of positive roots of $\Delta$ and $m_{\lambda}(\lambda)=1$.
[…]
(2.1) $\dim V(\lambda)=\sum_{\mu\in\Pi(\lambda)}m_{\lambda}(\mu)l(\mu)$
where $\Pi(\lambda):=\{\mu:\mu\mbox{ is a dominant weight such that }\mu\preceq\lambda\}$ and $l(\mu)$ denotes the length of the $\mathcal{W}$-orbit $\operatorname{Orb}_{\mathcal{W}}(\mu)$.
3. Multiplicities and dimensions for groups of type $A_{\ell}$
In order to be ready for computing certain multiplicities required along the section, let us start by giving a system of simple roots of type $A_{\ell}$ and the fundamental weights associated.
Let $\{\alpha_{1},\alpha_{2},\cdots,\alpha_{\ell}\}$ be the simple roots system of type $A_{\ell}$ given by
$\alpha_{i}:=e_{i}-e_{i+1},\quad\quad 1\leq i\leq\ell.$
We can normalize as follows: $(\alpha_{i},\alpha_{i})=1$, $(\alpha_{i},\alpha_{i+1})=-1/2$ and $(\alpha_{i},\alpha_{i+k})=0$, $k>1$. Then, the fundamental weights are given by
$\lambda_{i}=\frac{1}{\ell+1}[(\ell-i+1)\alpha_{1}+2(\ell-i+1)\alpha_{2}+i(\ell-i+1)\alpha_{i}+i(\ell-i)\alpha_{i+1}+\cdots+i\alpha_{\ell}].$
The following result will useful for computing the dimensions at hand.
Lemma 3.1.
For any $1\leq i,j\leq\ell$ we have that
$(\ell+1)^{2}(\lambda_{i},\lambda_{j})$ $=i(j-1)(\ell-j+1)^{2}\cdot(-1/2)+ij\sum_{k=1}^{\ell-j+1}k.$
Proof.
Without loss of generality suppose that $i\leq j$, namely $j=i+r$ with $r\geq 0$. Thus, we have
$\lambda_{i}$ $=(\ell-i+1)\alpha_{1}+\cdots+i(\ell-i+1)\alpha_{i}+\cdots+i(\ell-i-r+1)\alpha_{j}+i(\ell-i-r)\alpha_{j+1}+\cdots+i\alpha_{\ell}$
$\lambda_{j}$ $=(\ell-j+1)\alpha_{1}+\cdots+(\ell-j+1)\alpha_{i}+\cdots+j(\ell-j+1)\alpha_{j}+j(\ell-j)\alpha_{j+1}+\cdots+j\alpha_{j}.$
Then we can calculate
$(\ell+1)^{2}(\lambda_{i},\lambda_{j})$
$=(\ell-i+1)(\ell-j+1)+2(\ell-i+1)(\ell-j+1)\cdot(-1/2)$
$+2(\ell-i+1)(\ell-j+1)\cdot(-1/2)+4(\ell-i+1)(\ell-j+1)+6(\ell-i+1)(\ell-j+1)\cdot(-1/2)$
$+6(\ell-j+1)(\ell-i+1)\cdot(-1/2)+9(\ell-i+1)(\ell-j+1)+12(\ell-j+1)(\ell-i+1)\cdot(-1/2)$
$\vdots$
$+i(j-2)(\ell-j+1)(\ell-i-r+2)\cdot(-1/2)+i(j-1)(\ell-i-r+2)(\ell-j+1)+ij(\ell-j+1)(\ell-i-r+2)\cdot(-1/2)$
$+i(j-1)(\ell-i-r+1)(\ell-j+1)\cdot(-1/2)+ij(\ell-i-r+1)(\ell-j+1)+ij(\ell-i-r+1)(\ell-j)\cdot(-1/2)$
$+ij(\ell-i-r)(\ell-j+1)\cdot(-1/2)+ij(\ell-i-r)(\ell-j)+ij(\ell-i-r)(\ell-j-1)\cdot(-1/2)$
$+ij(\ell-i-r-1)(\ell-j)\cdot(-1/2)+ij(\ell-j-1)^{2}+ij(\ell-i-r-1)(\ell-j-2)\cdot(-1/2)$
$\vdots$
$+6ij\cdot(-1/2)+4ij+2ij\cdot(-1/2)$
$+2ij\cdot(-1/2)+ij$
$=i(j-1)(\ell-j+1)^{2}\cdot(-1/2)+ij(\ell-j+1)^{2}+ij(\ell-j+1)(\ell-j)\cdot(-1/2)$
$+ij(\ell-j)(\ell-j+1)\cdot(-1/2)+ij(\ell-j)^{2}+ij(\ell-j)(\ell-j-1)\cdot(-1/2)$
$+ij(\ell-j-1)(\ell-j)\cdot(-1/2)+ij(\ell-j-1)^{2}+ij(\ell-j-1)(\ell-j-2)\cdot(-1/2)$
$\vdots$
$+6ij\cdot(-1/2)+4ij+2ij\cdot(-1/2)$
$+2ij\cdot(-1/2)+ij$
$=i(j-1)(\ell-j+1)^{2}\cdot(-1/2)+ij\sum_{k=1}^{\ell-j+1}k^{2}-\sum_{k=1}^{\ell-j+1}k(k-1)$
$=i(j-1)(\ell-j+1)^{2}\cdot(-1/2)+ij\sum_{k=1}^{\ell-j+1}[k^{2}-k(k-1)]$
$=i(j-1)(\ell-j+1)^{2}\cdot(-1/2)+ij\sum_{k=1}^{\ell-j+1}k.$
Now we introduce a pivotal concept, specially in the next two sections.
Definition 3.2.
Let $G$ be a connected reductive simple algebraic group of type $A_{\ell}$ over $\overline{\mathbb{F}}_{p}$ of simply connected type and $\lambda$ be a dominant weight. We will say that $\lambda$ is admissible if the dimension of the $G(\overline{\mathbb{F}}_{p})$-module $V(\lambda)$ is at most $\ell^{4}/16$.
We see for instance.
Example 3.3.
When $G$ is as above with $\ell>2$, the dominant weight $\lambda_{1}+\lambda_{2}$ is admissible.
We need to compute the dimension of $V(\lambda_{1}+\lambda_{2})$ to verify the admissibility of $\lambda=\lambda_{1}+\lambda_{2}$. Observe first that $\lambda_{3}=\lambda-(\alpha_{1}+\alpha_{2})$, we then have to compute $m_{\lambda}(\lambda_{3}+\alpha_{1})$ and is idem for $m_{\lambda}(\lambda_{3}+\alpha_{2})$; we set $\mu=\lambda_{3}+\alpha_{2}=\lambda-\alpha_{1}$. Thus we have that
$m_{\lambda}(\mu)=2\cdot\frac{m_{\lambda}(\mu+\alpha_{2})(\mu+\alpha_{2},\alpha_{2})}{(\lambda+\delta,\lambda+\delta)-(\mu+\delta,\mu+\delta)}.$
Immediately we see that $m_{\lambda}(\mu+\alpha_{2})=m_{\lambda}(\lambda)=1$ and $(\mu+\alpha_{2},\alpha_{2})=1/2$. Thanks to Lemma 3.1 we deduce that
$(\lambda,\lambda)=\frac{5\ell-4}{2(\ell+1)}\quad\text{ and }\quad(\mu,\mu)=\frac{5\ell-4}{2(\ell+1)}.$
We then have
$m_{\lambda}(\mu_{2})$ $=2\cdot\frac{m_{\lambda}(\mu+\alpha_{2})(\mu+\alpha_{2},\alpha_{2})}{(\lambda,\lambda)-(\mu,\mu)+2(\lambda-\mu,\delta)}$
$=2\cdot\frac{1\cdot\frac{1}{2}}{\frac{5\ell-4}{2(\ell+1)}-\frac{5\ell-4}{2(\ell+1)}+1}$
$=1.$
Hence we compute the multiplicity of $\mu$ in $V(\lambda)$ as follows
$m_{\lambda}(\mu)$ $=\frac{(\mu+\alpha_{1},\alpha_{1})m_{\lambda}(\mu+\alpha_{1})+(\mu+\alpha_{2},\alpha_{2})m_{\lambda}(\mu+\alpha_{2})+(\mu+\alpha_{1}+\alpha_{2},\alpha_{1}+\alpha_{2})m_{\lambda}(\mu+\alpha_{1}+\alpha_{2})}{(\lambda+\delta,\lambda+\delta)-(\mu+\delta,\mu+\delta)}$
$=2\cdot\frac{1\cdot 1+1\cdot 1+1\cdot 1}{\frac{5\ell-4}{2(\ell+1)}-\frac{3(\ell-2)}{2(\ell+1)}+2}$
$=2.$
Thus, we conclude that
$\dim V(\lambda_{1}+\lambda_{2})=m_{\lambda}(\lambda)l(\lambda)+m_{\lambda}(\lambda_{3})l(\lambda_{3})$ $=\ell(\ell+1)+2\cdot\frac{(\ell-1)\ell(\ell+1)}{6}$
$=\ell(\ell+1)[1+(\ell-1)/3]$
$=\ell(\ell+1)(\ell+2)/3,$
as desired.
In contrast with example above we observe.
Example 3.4.
For all $\ell\geq 12$, the weight $\lambda=\lambda_{\ell-2}+\lambda_{\ell}$ is not admissible.
It suffices to show that the multiplicity of $\lambda_{\ell-3}\preceq\lambda$ is $3$ by resorting as follows
$\dim V(\lambda_{\ell-2}+\lambda_{\ell})\geq m_{\lambda}(\lambda_{\ell-3})l(\lambda_{\ell-3})$ $=3\cdot\frac{(\ell-2)(\ell-1)\ell(\ell+1)}{24}$
$=\frac{\ell^{4}(1-2/\ell)(1-1/\ell)(1+1/\ell)}{8}$
$>\frac{\ell^{4}}{16},$
by using $(1-2/\ell)(1-1/\ell)\geq(1-2/12)(1-1/12)>1/2$ for any integer $\ell\geq 12$ and $(1+1/\ell)\geq 1$. Thus, $\lambda$ cannot be admissible.
Then we shall compute $m_{\lambda}(\lambda_{\ell-3})$. We first note
$\mu:=\lambda_{\ell-3}=\lambda-(\alpha_{\ell}+\alpha_{\ell-1}+\alpha_{\ell-2})$
Thus, we have to compute
$m_{\lambda}(\mu)=\frac{(*)}{(\lambda,\lambda)-(\mu,\mu)+2(\lambda-\mu,\delta)},$
where
(*(*)( * ) $=2\sum_{\alpha\in\Phi^{+}}\sum_{i=1}^{\infty}(\mu+i\alpha,\alpha)m_{\lambda}(\mu+i\alpha)$
$=(\mu+(\alpha_{\ell}+\alpha_{\ell-1}+\alpha_{\ell-2}),\alpha_{\ell}+\alpha_{\ell-1}+\alpha_{\ell-2})m_{\lambda}(\mu+(\alpha_{\ell}+\alpha_{\ell-1}+\alpha_{\ell-2}))$
$+(\mu+(\alpha_{\ell}+\alpha_{\ell-1}),\alpha_{\ell}+\alpha_{\ell-1})m_{\lambda}(\mu+(\alpha_{\ell}+\alpha_{\ell-1}))$
$+(\mu+(\alpha_{\ell-1}+\alpha_{\ell-2}),\alpha_{\ell-1}+\alpha_{\ell-2})m_{\lambda}(\mu+(\alpha_{\ell-1}+\alpha_{\ell-2}))$
$+(\mu+\alpha_{\ell},\alpha_{\ell})m_{\lambda}(\mu+\alpha_{\ell})+(\mu+\alpha_{\ell-1},\alpha_{\ell-1})m_{\lambda}(\mu+\alpha_{\ell-1})$
$+(\mu+\alpha_{\ell-2},\alpha_{\ell-2})m_{\lambda}(\mu+\alpha_{\ell-2}).$
By computing -term by term- as in Example 3.3 we may infer that
$m_{\lambda}(\mu+\alpha_{\ell})(\mu+\alpha_{\ell},\alpha_{\ell})$ $=1\cdot 2$
$m_{\lambda}(\mu+\alpha_{\ell-1})(\mu+\alpha_{\ell-1},\alpha_{\ell-1})$ $=1\cdot 2$
$m_{\lambda}(\mu+\alpha_{\ell-2})(\mu+\alpha_{\ell-2},\alpha_{\ell-2})$ $=1\cdot 2$
$m_{\lambda}(\mu+\alpha_{\ell-1}+\alpha_{\ell})(\mu+\alpha_{\ell-1}+\alpha_{\ell},\alpha_{\ell-1}+\alpha_{\ell})$ $=1\cdot 2$
$m_{\lambda}(\mu+\alpha_{\ell-1}+\alpha_{\ell-2})(\mu+\alpha_{\ell-1}+\alpha_{\ell-2},\alpha_{\ell-1}+\alpha_{\ell-2})$ $=1\cdot 2$
$m_{\lambda}(\mu+\alpha_{\ell-2}+\alpha_{\ell-1}+\alpha_{\ell})(\mu+\alpha_{\ell-2}+\alpha_{\ell-1}+\alpha_{\ell},+\alpha_{\ell-2}+\alpha_{\ell-1}+\alpha_{\ell})$ $=1\cdot 2.$
Finally, since
$(\lambda,\lambda)-(\mu,\mu)+2(\lambda-\mu,\delta)=(\lambda,\lambda)-(\mu,\mu)+2(\alpha_{\ell}+\alpha_{\ell-1}+\alpha_{\ell-2},\delta)=6-4+2\cdot 3=8,$
we have:
$m_{\lambda}(\mu)=2\cdot\frac{2\cdot 1+2\cdot 1+2\cdot 1+2\cdot 1+2\cdot 1+2\cdot 1}{8}=2\cdot\frac{12}{8}=3.$
Hence $\lambda=\lambda_{\ell-2}+\lambda_{\ell}$ is not admissible.
Remark 3.5.
It is worth to emphasizing the obvious inequality
$\dim V(\lambda)\geq m_{\lambda}(\mu)l(\mu)\geq l(\mu)$
that we will frequently use for any dominant weight $\mu\prec\lambda$.
Proposition 3.6.
Let $G$ be as above with $\ell\geq 15$. Then, all weights in the following table are admissible.
$\lambda$ dimension of $V(\lambda)$
00 $1$
$\lambda_{1}$, $\lambda_{\ell}$ $\ell+1$
$\lambda_{2}$, $\lambda_{\ell-1}$ $\ell(\ell+1)/2$
$\lambda_{3}$, $\lambda_{\ell-2}$ $(\ell-1)\ell(\ell+1)/6$
$\lambda_{4}$, $\lambda_{\ell-3}$ $(\ell-2)(\ell-1)\ell(\ell+1)/24$
$2\lambda_{1}$, $2\lambda_{\ell}$ $(\ell+2)(\ell+1)/2$
$3\lambda_{1}$, $3\lambda_{\ell}$ $(\ell+3)(\ell+2)(\ell+1)/6$
$\lambda_{1}+\lambda_{\ell}$ $\ell(\ell+2)$
$\lambda_{1}+\lambda_{\ell-1}$, $\lambda_{2}+\lambda_{\ell}$ $(\ell+2)(\ell+1)(\ell-1)/2$
$\lambda_{1}+2\lambda_{\ell}$, $2\lambda_{1}+\lambda_{\ell}$ $(\ell+1)\ell(\ell+3)/2$
$\lambda_{1}+\lambda_{2}$, $\lambda_{\ell-1}+\lambda_{\ell}$ $\ell(\ell+1)(\ell+2)/3$
In addition for $\ell\geq 16$, the weights $4\lambda_{1}$ and $4\lambda_{\ell}$ are also admissible whose associated irreducible representations have dimension $(\ell+4)(\ell+3)(\ell+2)(\ell+1)/24$.
Proof.
Firstly, we shall prove that, for $\ell\geq 15$, the following weights
[…]
are admissible. We start remarking that $V(k\lambda_{\ell})$ is irreducible by definition and is contained in the symmetric power tensor product $\operatorname{Sym}^{k}(V)$, where $V$ is the standard irreducible representation of dimension $\ell+1$. Now, let us start by deducing the dimension of the weight space $V(3\lambda_{\ell})$. Note that the dominant weights $\mu\preceq 3\lambda_{\ell}$ are: $3\lambda_{\ell}$, $\lambda_{\ell-1}+\lambda_{\ell}$, and $\lambda_{\ell-2}$. In addition, their stabilizers are contained in the reflection groups of type $A_{\ell-1}$, $A_{\ell-2}$ and $A_{\ell-3}\times A_{2}$, respectively. Since its multiplicities are at least $1$, by formula (2.1) we can observe
$\dim V(3\lambda_{\ell})$ $\geq l(3\lambda_{\ell})+l(\lambda_{\ell-1}+\lambda_{\ell})+l(\lambda_{\ell-2})$
$=(\ell+1)+\ell(\ell+1)+\ell(\ell-1)(\ell+1)/6$
$=(\ell+1)(\ell+2)(\ell+3)/6$
$=\dim\operatorname{Sym}^{3}(V).$
This then implies that $V(3\lambda_{\ell})=\operatorname{Sym}^{3}(V)$ and $\dim V(3\lambda_{\ell})=(\ell+1)(\ell+2)(\ell+3)/6\leq\ell^{4}/6$. By duality we infer the dimensions of $V(3\lambda_{1})$. Similarly, we can deduce that the anti-symmetric power tensor product $\Lambda^{i+1}(V)=V(\lambda_{\ell-i})$ for $i=2,3$ by observing that $V(\lambda_{\ell-i})\subseteq\Lambda^{i+1}(V)$ and
$\dim V(\lambda_{\ell-i})\geq l(\lambda_{\ell-i})=\frac{(\ell+1)!}{(\ell-i)!(i+1)!}=\binom{\ell+1}{i+1}=\dim\Lambda^{i+1}(V).$
[…]
Now, when $\ell\geq 16$ we may check that $\lambda=4\lambda_{\ell}$ is admissible by verifying first that the dominant weights $\mu\preceq 4\lambda_{\ell}$ are: $4\lambda_{\ell}$, $\lambda_{\ell-1}+2\lambda_{\ell}$, $2\lambda_{\ell-1}$, $\lambda_{\ell-2}+\lambda_{\ell-1}$ and $\lambda_{\ell-3}$, and later to check
$\dim V(4\lambda_{\ell})\geq l(4\lambda_{\ell})+l(\lambda_{\ell-1}+2\lambda_{\ell})+l(2\lambda_{\ell-1})+l(\lambda_{\ell-2}+\lambda_{\ell})+l(\lambda_{\ell-3})=\dim\operatorname{Sym}^{4}(V).$
This implies that $V(4\lambda_{\ell})=\operatorname{Sym}^{4}(V)$ and $\dim V(4\lambda_{\ell})=(\ell+1)(\ell+2)(\ell+3)(\ell+4)/24\leq\ell^{4}/16$. Then, by duality, we deduce the dimension of $V(4\lambda_{1})$. Finally, we may note that the remaining weights were studied by Lübeck in [11, §5]. ∎
Remark 3.7.
Notice that the argument used to prove $\Lambda^{i+1}(V)=V(\lambda_{\ell-i})$ for $i=2,3$ (resp. $V(i\lambda_{\ell})=\operatorname{Sym}^{i}(V)$ for $i=3,4$) works for any integer $i$ such that $11\leq i\leq\lfloor\ell/2\rfloor+$ 1. This tell us in particular that $\Lambda^{i+1}(V)$ (resp. $\operatorname{Sym}^{i}(V)$) is irreducible. Moreover, by duality we may observe that
$\dim V(\lambda_{i+1})=\dim\Lambda^{i+1}(V),\quad(\mbox{resp. }\dim V(i\lambda_{1})=\dim\operatorname{Sym}^{i}(V)).$
4. Representations of finite groups of type $A_{\ell}$
In order to determine all admissible weights when $G$ is a finite group of Lie type $A_{\ell}$ over $\mathbb{\overline{F}}_{p}$ we shall proceed by a series of steps to reduce the possibilities according to whether the coefficients are zeroes or not. Let $\lambda=a_{1}\lambda_{1}+a_{2}\lambda_{2}+\cdots+a_{\ell}\lambda_{\ell}$ be a dominant weight. The first step, nullity of coefficients, consists of identifying all indexes $i$ such that $a_{i}=0$. Each of the other reductions will be associated to a nonzero coefficient $a_{i}$, where we determine all the possibilities to the other nonzero coefficients $a_{j}$ reducing us exclusively to admissible weights. Explicitly, we start with.
Nullity of coefficients -- Claim: $a_{i}\neq 0$ only if $i\leq 4$ or $i\geq\ell-3$.
Suppose $a_{i}\neq 0$ for some $i$ satisfying $45\leq i\leq\ell-$ 4. This implies that the $\mathcal{W}$-stabilizer of $\lambda$ is contained in a reflection group of type $A_{i-1}\times A_{\ell-i}$. Thus,
$\dim V(\lambda)\geq l(\lambda)=\frac{(\ell+1)!}{i!(\ell+1-i)!}.$
We shall prove that $(\ell+1)!/i!(\ell+1-i)!>\ell^{4}/16$, for any $45\leq i\leq\ell-$ 4. To do this, assume firstly that $i\leq\ell+1-i$. In this case, we need to show that
$\frac{(\ell+2-i)(\ell+3-i)\cdots(\ell-3)(\ell-2)(\ell-1)\ell(\ell+1)}{i!}>\frac{\ell^{4}}{16}.$
We shall first verify the inequality for $i=5$, that is,
$\frac{(\ell-3)(\ell-2)(\ell-1)\ell(\ell+1)}{5!}>\frac{\ell^{4}}{16}$ $\iff\frac{(\ell-3)(\ell-2)(\ell-1)(\ell+1)}{3\cdot 5}>\frac{\ell^{3}}{2}.$
In fact, since $\ell\geq 15$, we can infer that $(\ell+1)/15>1$ and $151-j/\ell\geq 1-j/$ 15 for $1\leq j\leq 3$. Then, by combining these facts we can deduce
$\frac{(\ell-3)(\ell-2)(\ell-1)(\ell+1)}{3\cdot 5}$ $>(\ell-3)(\ell-2)(\ell-1)$
$\geq\ell^{3}(1-3/\ell)(1-2/\ell)(1-1/\ell)$
$\geq\ell^{3}(1-3/15)(1-2/15)(1-1/15)$
$\geq\ell^{3}\frac{12}{15}\cdot\frac{13}{15}\cdot\frac{14}{15}$
$>\frac{\ell^{3}}{2}.$
Thus, the case $i=5$ follows by multiplying by $\ell/8$.
The general case $i\leq\ell+1-i$ is then deduced from the base case $i=5$ by noting that
$\frac{(\ell+2-i)(\ell+3-i)\cdots(\ell-4)(\ell-3)(\ell-2)(\ell-1)\ell(\ell+1)}{i(i-1)\cdots 6\cdot 5!}$
$=\left(\frac{(\ell+2-i)(\ell+2-(i+1))\cdots(\ell+2-6)}{i(i-1)\cdots 6}\right)\left(\frac{(\ell-3)(\ell-2)(\ell-1)\ell(\ell+1)}{5!}\right)$
and the first parenthesis is greater than 1 by using the condition $\ell\geq 15$. Finally, the case $\ell+1-i<i\leq\ell$ follows by using a symmetric argument. Therefore, we reduces us to
$\lambda=a_{1}\lambda_{1}+a_{2}\lambda_{2}+a_{3}\lambda_{3}+a_{4}\lambda_{4}+a_{\ell-3}\lambda_{\ell-3}+a_{\ell-2}\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}.$
Liberty of coefficients -- Assuming that $a_{\ell-3}\neq 0$ we shall show that $a_{\ell-3}=1$ and $a_{i}=0$ for any $i\leq 4$ and $i\geq\ell-2$. In fact, if $a_{\ell-j}\neq 0$ for all $j=0,1,2$ then the $\mathcal{W}$-stabilizer of $\lambda$ is contained in a reflection group of type $A_{\ell-4}$. We then see that
(4.1) $\dim V(\lambda)\geq l(\lambda)=\frac{(\ell+1)!}{(\ell-3)!}=(\ell-2)(\ell-1)\ell(\ell+1)>\frac{\ell^{4}}{16}.$
On the other hand, when exactly two coefficients $a_{\ell-j}$ are nonzero, the $\mathcal{W}$-stabilizer of $\lambda$ is contained in a reflection group of type $A_{\ell-3}\times A_{2}$. An argument as above shows that $\dim V(\lambda)>\ell^{4}/16$. Next, if just one coefficient $a_{\ell-j}$ is nonzero then the $\mathcal{W}$-stabilizer of $\lambda$ is contained in a reflection group of type $A_{\ell-4}\times A_{1}\times A_{1}$ or $A_{\ell-4}\times A_{2}$ according to $j=1$ or not. Thus, we may rule out this possibility of $\lambda$. It follows that $a_{\ell-j}=0$ for all $j=0,1,2$ and by a symmetry argument one checks that $a_{i}=0$ for all $i=1,2,3$. Thus, we reduce to $\lambda=a_{\ell-3}\lambda_{\ell-3}$.
We now will prove that $a_{\ell-3}=1$ by supposing the contrary i.e., $a_{\ell-3}\geq 2$. Then, we observe that the $(\ell-4)$-coefficient of
$\lambda-\alpha_{\ell-3}=a_{\ell-3}\lambda_{\ell-3}-(-\lambda_{\ell-4}+2\lambda_{\ell-3}-\lambda_{\ell-2})=\lambda_{\ell-4}+(a_{\ell-3}-2)\lambda_{\ell-3}+\lambda_{\ell-2}$
is one, which is not possible by Nullity of Coefficients. Whence $a_{\ell-3}=1$ and then $\lambda=\lambda_{\ell-3}$ is admissible by Proposition 3.6.
Now assume that $a_{\ell-3}=0$ and $a_{\ell-2}\neq 0$, that is,
$\lambda=a_{1}\lambda_{1}+a_{2}\lambda_{2}+a_{3}\lambda_{3}+a_{4}\lambda_{4}+a_{\ell-2}\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}.$
We shall prove that $a_{i}=0$ for $1\leq i\leq 4$. Suppose first $a_{1}\neq 0$ and $a_{2}=a_{3}=a_{4}=0$. We note that the $\mathcal{W}$-stabilizer of $\lambda$ is contained in a reflection group of type $A_{\ell-4}$ or $A_{\ell-4}\times A_{1}$ or $A_{\ell-4}\times A_{2}$ according to $a_{\ell-1}\neq 0\neq a_{\ell}$ or not. Then, reasoning as above, we can deduce that $\dim V(\lambda)>\ell^{4}/16$. Similarly each possibility of some $a_{i}\neq 0$ for $1\leq i\leq 4$ discards $\lambda$ as admissible weight, and thus
$\lambda=a_{\ell-2}\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}.$
We now show that $a_{\ell-2}\leq 1$. Otherwise, assuming that $a_{\ell-2}\geq 2$ one observes that the coefficients at $(\ell-3)$ and $(\ell-1)$ of the dominant weight $\lambda-a_{\ell-2}$ are nonzero, but this is not possible by reduction above. Hence, we reduce our analysis to study
$\lambda=\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}\quad\text{or }\quad\lambda=a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}.$
When $a_{\ell-2}=1$ and $a_{\ell-1}\geq 2$ then the term $2\lambda_{\ell-2}$ appears in $\lambda-\alpha_{\ell-1}$, which is impossible by applying the reasoning above. So, any $\lambda$ of the form $\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$ with $a_{\ell-1}\geq 2$ is not admissible and we reduce us to $\lambda$ of the form $\lambda=\lambda_{\ell-2}+\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$ or $\lambda=\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}$. For the case $\lambda=\lambda_{\ell-2}+\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$, if $a_{\ell}\geq 2$ we may observe similarly that the term $2\lambda_{\ell-1}$ appears in the dominant weight $\lambda-\alpha_{\ell-1}$ which is a contradiction by iterating the previous reasoning. This reduces our analysis to study the dominant weights $\lambda_{\ell-2}+\lambda_{\ell-1}+\lambda_{\ell}$ and $\lambda_{\ell-2}+\lambda_{\ell-1}$. In the same vein, we discard the possibility of a dominant weight of the form $\lambda_{\ell-2}+a_{\ell}\lambda_{\ell}$ with $a_{\ell}\geq 4$. The cases $a_{\ell}=2,3$ are reduced again to analyze the dominant weights $\lambda_{\ell-2}+\lambda_{\ell-1}+\lambda_{\ell}$ and $\lambda_{\ell-2}+\lambda_{\ell-1}$. Finally, when $\lambda=\lambda_{\ell-2}+a_{\ell}\lambda_{\ell}$ with $a_{\ell}\geq 2$ cannot be admissible by using similar arguments as above. Therefore, we only need to determinate the dimensions of $V(\lambda)$ for the following dominants weights
$\lambda_{\ell-2}+\lambda_{\ell-1},\quad\lambda_{\ell-2}+\lambda_{\ell-1}+\lambda_{\ell},\quad\lambda_{\ell-2}+\lambda_{\ell},\quad\lambda_{\ell-2}.$
In the first case we see that $\lambda_{\ell-3}+\lambda_{\ell}\prec\lambda_{\ell-2}+\lambda_{\ell-1}$. Thus, we can deduce as before that
$\dim V(\lambda_{\ell-2}+\lambda_{\ell-1})\geq l(\lambda_{\ell-3}+\lambda_{\ell})$ $=\frac{\ell^{4}(1-2/\ell)(1-1/\ell)(1+1/\ell)}{6}>\frac{\ell^{4}(1-2/\ell)(1-1/\ell)}{6}>\ell^{4}/12>\ell^{4}/16,$
since $(1-2/\ell)(1-1/\ell)\geq((1-2/15)(1-1/15)>1/2$ for any $\ell\geq 15$. In the same way, we can rule out the second case by using that $\lambda_{\ell-3}+2\lambda_{\ell}\prec\lambda_{\ell-2}+\lambda_{\ell-1}+\lambda_{\ell}$. Note that $\lambda_{\ell-2}+\lambda_{\ell}$ is not admissible by appealing to Example 3.4 and $\lambda_{\ell-2}$ is admissible by Proposition 3.6.
Next, suppose $a_{\ell-3}=a_{\ell-2}=0$ and $a_{\ell-1}\neq 0$:
$\lambda=a_{1}\lambda_{1}+a_{2}\lambda_{2}+a_{3}\lambda_{3}+a_{4}\lambda_{4}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}.$
If $a_{\ell-1}\geq 2$ we note again that $\lambda_{\ell-2}$ and $\lambda_{\ell}$ appear in the writing of the dominant weight $\lambda-\alpha_{\ell-1}$, which is impossible as was showed previously. Thus, we reduce to $a_{\ell-1}=1$, i.e.,
$\lambda=a_{1}\lambda_{1}+a_{2}\lambda_{2}+a_{3}\lambda_{3}+a_{4}\lambda_{4}+\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}.$
In this case we may prove that $a_{i}=0$ for $2\leq i\leq 4$. Here, many cases arise for the $\mathcal{W}$-stabilizer of $\lambda$ according to the vanished for some of these $a_{i}$ ($2\leq i\leq 4)$. For instance, if $a_{2}\neq 0$ and $a_{3}=a_{4}=0$ we see that $\mathcal{W}$-stabilizer of $\lambda$ is contained in a reflection group of type $A_{1}\times A_{\ell-4}$ or $A_{\ell-4}$ according to $a_{1}=0$ or not. Thus, we may verify that $\dim V(\lambda)>\ell^{4}/16$, i.e., $\lambda$ is not admissible. Indeed, in any case we may deduce that the dimension $V(\lambda)$ is bigger than the bound $\ell^{4}/16$. This implies
$\lambda=a_{1}\lambda_{1}+\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}.$
Reasoning as above one verifies that $a_{1}$ and $a_{\ell}$ cannot be simultaneously nonzero. In fact, assume first that $a_{1}\geq 2$ we obtain the dominant weight
$\lambda-\alpha_{1}=(a_{1}-2)\lambda_{1}+\lambda_{2}+\lambda_{\ell-1}+a_{\ell}\lambda_{\ell},$
which is ruled out by computations as above. On the other hand, if $a_{\ell}\geq 2$ then $2\lambda_{\ell-1}$ appears in the writing of $\lambda-\alpha_{\ell}$, which is impossible too. Moreover, we may discard $\lambda=\lambda_{1}+\lambda_{\ell-1}+\lambda_{\ell}$ by arguing as in Example 3.4. This lefts us the dominant weights $\lambda_{1}+\lambda_{\ell-1}$, $\lambda_{\ell-1}+\lambda_{\ell}$ and $\lambda_{\ell-1}$ which are admissible by Proposition 3.6.
Finally suppose that $a_{\ell-3}=a_{\ell-2}=a_{\ell-1}=0$ and $a_{\ell}\neq 0$, i.e., we will deal with
$\lambda=a_{1}\lambda_{1}+a_{2}\lambda_{2}+a_{3}\lambda_{3}+a_{4}\lambda_{4}+a_{\ell}\lambda_{\ell}$
We may verify, similarly as above, that $a_{3}=a_{4}=0$, reducing our analysis to
$\lambda=a_{1}\lambda_{1}+a_{2}\lambda_{2}+a_{\ell}\lambda_{\ell}.$
Observe that if $a_{2}\neq 0$ then $a_{1}=0$ by using a dual reasoning to the previous case. This implies that $\lambda=\lambda_{2}+a_{\ell}\lambda_{\ell}$, which gives just one new admissible weight $\lambda=\lambda_{2}+\lambda_{\ell}$. Finally, if $a_{2}=0$ then we have that $\lambda=a_{1}\lambda_{1}+a_{\ell}\lambda_{\ell}$ and we get the new admissible weights $\lambda_{1}+\lambda_{\ell}$, $2\lambda_{1}+\lambda_{\ell}$, $\lambda_{1}+2\lambda_{\ell}$ and $a_{\ell}\lambda_{\ell}$ for $a_{\ell}=1,2,3,4$.
In summary, we have proved the first of our main results:
Theorem 4.1.
Let $G$ be a finite group of Lie type $A_{\ell}$ over $\mathbb{\overline{F}}_{p}$. If $\ell\geq 15$ then the list giving in Proposition 3.6 exhausts the admissible weights.
5. Multiplicities and dimension in finite classical groups
In this section we deal with the rest of finite classical groups. Let $G$ be a connected reductive simple algebraic group of type $B_{\ell}$, $C_{\ell}$ or $D_{\ell}$ of simply connected type. By analogy to Definition 3.2 we present the notion of admissibility for these types.
Definition 5.1.
Let $G$ be a connected reductive simple algebraic group of type $B_{\ell}$, $C_{\ell}$ or $D_{\ell}$ over $\overline{\mathbb{F}}_{p}$ of simply connected type and $\lambda$ be a dominant weight. We will say that $\lambda$ is admissible if the dimension of the $G(\overline{\mathbb{F}}_{p})$-module $V(\lambda)$ is at most $\ell^{4}$.
We shall begin the desired classification of admissible weights by considering $G$ of type $B_{\ell}$.
5.1. Type $B_{\ell}$
In order to compute the multiplicities involved in this section, we start by considering $\{\alpha_{1},\alpha_{2},\cdots,\alpha_{\ell}\}$ the simple roots system (of type $B_{\ell}$) given by
$\alpha_{1}:=e_{\ell}\quad\text{and}\quad\alpha_{i}:=e_{\ell-i+1}-e_{\ell-i+2},\quad\quad 2\leq i\leq\ell,$
where $\{e_{1},e_{2},\cdots,e_{\ell}\}$ is an orthogonal basis for $E\cong\mathbb{R}^{\ell}$. We then note that $(\alpha_{1},\alpha_{1})=1$, $(\alpha_{1},\alpha_{2})=-1$, and for $i,j>1$ with $i\neq j$ we see that
(5.1) $(\alpha_{i},\alpha_{j})=\begin{cases}\phantom{-}2,&\text{ if }i=j\\ -1,&\text{ if }i=j\pm 1\\ \phantom{-}0,&\text{ otherwise }.\\ \end{cases}$
Then the fundamental weights $\lambda_{i}$ ($1\leq i\leq\ell$), such that $(\lambda_{i},\alpha_{j})=\delta_{ij}$, are given by
$\lambda_{1}$ $=(1/2)(\alpha_{\ell}+2\alpha_{\ell-1}+\cdots+\ell\alpha_{1})$
$\lambda_{i}$ $=\alpha_{\ell}+2\alpha_{\ell-1}+\cdots+(\ell-i)\alpha_{i+1}+(\ell-i+1)(\alpha_{i}+\alpha_{i-1}+\cdots+\alpha_{2}+\alpha_{1}),\quad 1<i<\ell$
$\lambda_{\ell}$ $=\alpha_{\ell}+\alpha_{\ell-1}+\cdots+\alpha_{2}+\alpha_{1}.$
In particular, we have that $(\lambda_{i},\lambda_{i})=\ell-i+1.$ In addition, the half sum of the positive roots is
$\delta=\frac{1}{2}\sum_{i=1}^{\ell}(\ell-i+1)(\ell+i-1)\alpha_{i}.$
It is easy to verify
$(\alpha_{i},\delta)=\begin{cases}\frac{1}{2},&\text{ if }i=1\\ 1,&\text{ otherwise }.\\ \end{cases}$
Let us start with the first example of admissibility for this type.
Example 5.2.
For all $\ell\geq 12$, the weight $\lambda=4\lambda_{\ell}$ is admissible. To see that, it suffices to show that $\dim\operatorname{Sym}^{4}(V)$ is less or equal to $\ell^{4}$ since $V(4\lambda_{\ell})\subseteq\operatorname{Sym}^{4}(V)$. In fact, we observe that
$\dim\operatorname{Sym}^{4}(V)$ $=\frac{(2\ell+1)(2\ell+2)(2\ell+3)(2\ell+4)}{24}$
$=\frac{(2\ell+1)(\ell+1)(2\ell+3)(\ell+2)}{6}$
$=\ell^{4}\cdot\frac{(2+1/\ell)(1+1/\ell)(2+3/\ell)(1+2/\ell)}{6}$
$\leq\ell^{4},$
by verifying the inequality
$(2+1/\ell)(1+1/\ell)(2+3/\ell)(1+2/\ell)\\ \leq(2+1/12)(1+1/12)(2+3/12)(1+2/12)\leq 6,$
for any $\ell\geq 12$. Thus $\lambda=4\lambda_{\ell}$ is an admissible weight.
To find all the admissible weights $\lambda$ and compute dimensions of certain $G$-modules $V(\lambda)$, we will need the following lemmas.
Lemma 5.3.
For all $\ell\geq 12$, the multiplicities of all dominant weights $\mu\preceq i\lambda_{\ell}$, $i=3,4$, are given in the following table:
highest weight dominant weight $\mu\preceq\lambda$ multiplicity of $\mu$
$3\lambda_{\ell}$ $3\lambda_{\ell}$ $1$
$\lambda_{\ell-1}+\lambda_{\ell}$ $1$
$\lambda_{\ell-2}$ $1$
$2\lambda_{\ell}$ $1$
$\lambda_{\ell-1}$ $1$
[…]
00 $\ell$
$4\lambda_{\ell}$ $4\lambda_{\ell}$ $1$
$\lambda_{\ell-1}+2\lambda_{\ell}$ $1$
$2\lambda_{\ell-1}$ $1$
$\lambda_{\ell-2}+\lambda_{\ell}$ $1$
$\lambda_{\ell-3}$ $1$
$3\lambda_{\ell}$ $1$
$\lambda_{\ell-1}+\lambda_{\ell}$ $1$
$\lambda_{\ell-2}$ $1$
[…]
00 $\ell(\ell-1)/2$
Proof.
Let $\lambda=3\lambda_{\ell}$. We shall compute those multiplicities $m_{\lambda}(\mu)$ where $2\lambda_{\ell}\prec\mu\preceq 3\lambda_{\ell}$. Since $\mu:=\lambda_{\ell-1}+\lambda_{\ell}=\lambda-\alpha_{\ell}$ we have that $(\lambda,\lambda)=9$ and $(\mu,\mu)=5$. In addition as $(\alpha_{\ell},\delta)=1$ and $(\lambda,\alpha_{\ell})=3(\lambda_{\ell},\alpha_{\ell})=1$, we have
$m_{\lambda}(\mu)=2\cdot\frac{m_{\lambda}(\mu+\alpha_{\ell})(\mu+\alpha_{\ell},\alpha_{\ell})}{(\lambda,\lambda)-(\mu,\mu)+2(\alpha_{\ell},\delta)}=2\cdot\frac{1\cdot 3}{9-5+2\cdot 1}=1.$
Next, we compute the multiplicity of $\lambda_{\ell-2}=\lambda-(\alpha_{\ell-1}+2\alpha_{\ell})$. To do this, we observe that the positive roots involved in Fruedenthal formula are $\alpha_{\ell-1}$, $\alpha_{\ell}$, $2\alpha_{\ell}$ and $\alpha_{\ell-1}+\alpha_{\ell}$. Notice, at this point, that $\lambda$ is not included in the formula since $\mu+\alpha_{\ell-1}+2\alpha_{\ell}$ is not of the form $\mu+i\alpha$ for some positive root $\alpha$ by appealing to [3, Plate II].
Set $\mu:=\lambda-(\alpha_{\ell-1}+\alpha_{\ell})$.
  1. (1)
    Set $\mu_{\ell-1}:=\mu+\alpha_{\ell-1}=\lambda-\alpha_{\ell}$. We first see that $m_{\lambda}(\mu_{\ell-1}+\alpha_{\ell})=1$ and $(\mu_{\ell}+\alpha_{\ell},\alpha_{\ell})=9$. In addition, we already have $(\lambda,\lambda)=9$, and we verify that $(\mu_{\ell-1},\mu_{\ell-1})=5$. Thus, we conclude
    $m_{\lambda}(\mu_{\ell-1})=2\cdot\frac{m_{\lambda}(\mu_{\ell-1}+\alpha_{\ell})(\mu_{\ell-1}+\alpha_{\ell},\alpha_{\ell})}{(\lambda,\lambda)-(\mu_{\ell-1},\mu_{\ell-1})+2(\alpha_{\ell},\delta)}=2\cdot\frac{1\cdot 3}{9-5+2}=1.$
  2. (2)
    For $\mu_{\ell}:=\mu+\alpha_{\ell}=\lambda-\alpha_{\ell-1}$, we can conclude that $m_{\lambda}(\mu_{\ell})=0$ since $(\mu_{\ell}+\alpha_{\ell-1},\alpha_{\ell-1})=0$.
Hence, by verifying that $(\mu+(\alpha_{\ell-1}+\alpha_{\ell}),\alpha_{\ell-1}+\alpha_{\ell})=3$, $(\mu,\mu)=5$ and $(\alpha_{\ell-1}+\alpha_{\ell},\delta)=2$
$m_{\lambda}(\mu)$ $=2\cdot\frac{m_{\lambda}(\mu_{\ell-1}+\alpha_{\ell-1})(\mu_{\ell-1}+\alpha_{\ell-1},\alpha_{\ell-1})+m_{\lambda}(\mu_{\ell}+\alpha_{\ell})(\mu_{\ell}+\alpha_{\ell},\alpha_{\ell})+m_{\lambda}(\lambda)(\lambda_{\ell},\alpha_{\ell-1}+\alpha_{\ell})}{(\lambda,\lambda)-(\lambda_{\ell-2},\lambda_{\ell-2})+2(\alpha_{\ell-1}+\alpha_{\ell},\delta)}=1$
$=2\cdot\frac{1\cdot 1+0+1\cdot 3}{9-5+4}=1.$
We continue with the computing of $m_{\lambda}(\lambda-2\alpha_{\ell})$. To this end we may deduce as above that $m_{\lambda}(\lambda-\alpha_{\ell})=1$ and next we shall compute the multiplicity $m_{\lambda}(\mu+2\alpha_{\ell})$. We can verify that $(\mu+\alpha_{\ell},\alpha_{\ell})=1$, $(\mu+2\alpha_{\ell},\alpha_{\ell})=3$ and $(\mu,\mu)=5$. Thus, we have
$m_{\lambda}(\lambda-2\alpha_{\ell})$ $=2\cdot\frac{m_{\lambda}(\lambda-\alpha_{\ell})(\lambda-\alpha_{\ell},\alpha_{\ell})+m_{\lambda}(\lambda)(\lambda,\alpha_{\ell})}{(\lambda,\lambda)-(\mu,\mu)+2(2\alpha_{\ell},\delta)}$
$=2\cdot\frac{1\cdot 1+1\cdot 3}{9-5+4}=1.$
Going back to computing of $m_{\lambda}(\mu)$ with $\mu=\lambda_{\ell-2}=\lambda-(\alpha_{\ell-1}+2\alpha_{\ell})$ we have
$m_{\lambda}(\mu)=2\cdot\frac{(*)}{(\lambda,\lambda)-(\mu,\mu)+2(\alpha_{\ell-1}+2\alpha_{\ell},\delta)}$
where
(*(*)( * ) $=(\mu+\alpha_{\ell},\alpha_{\ell})m_{\lambda}(\mu+\alpha_{\ell})+(\mu+\alpha_{\ell-1})m_{\lambda}(\mu+\alpha_{\ell-1},\alpha_{\ell-1})$
$+(\mu+2\alpha_{\ell},\alpha_{\ell})m_{\lambda}(\mu+2\alpha_{\ell})+(\mu+\alpha_{\ell-1}+\alpha_{\ell},\alpha_{\ell-1}+\alpha_{\ell})m_{\lambda}(\mu+\alpha_{\ell-1}+\alpha_{\ell})$
$=(\mu+\alpha_{\ell},\alpha_{\ell})m_{\lambda}(\lambda-(\alpha_{\ell-1}+\alpha_{\ell}))+(\mu+\alpha_{\ell-1},\alpha_{\ell-1})m_{\lambda}(\lambda-2\alpha_{\ell})$
$+(\mu+2\alpha_{\ell},\alpha_{\ell})m_{\lambda}(\lambda-\alpha_{\ell-1})+(\mu+\alpha_{\ell-1}+\alpha_{\ell},\alpha_{\ell-1}+\alpha_{\ell})m_{\lambda}(\lambda-\alpha_{\ell}).$
Computing as above we see that
$(\lambda,\lambda)-(\mu,\mu)+2(\alpha_{\ell-1}+2\alpha_{\ell},\delta)=9-3+6=12.$
Now, by using the multiplicities calculated above and computing the inner products, we have
$m_{\lambda}(\mu)=\frac{2[1\cdot 2+1\cdot 2+0\cdot 4+1\cdot 2]}{12}=1.$
The remaining multiplicities $m_{\lambda}(\mu)$, with $\mu\preceq 2\lambda$ were compute in [11].
Finally, the case $\lambda=4\lambda_{\ell}$ is longer but completely similar than the one above by computing the multiplicity of $\lambda_{\ell-1}+2\lambda_{\ell}$ as first step. ∎
Lemma 5.4.
For all $\ell\geq 12$, the multiplicities of all dominant weights $\mu\preceq\lambda_{\ell-i}$, $i=2,3$, are given in the following table:
highest weight dominant weight $\mu\preceq\lambda$ multiplicity of $\mu$
$\lambda_{\ell-2}$ $\lambda_{\ell-2}$ $1$
$\lambda_{\ell-1}$ $1$
$\lambda_{\ell}$ $\ell-1$
00 $\ell$
$\lambda_{\ell-3}$ $\lambda_{\ell-3}$ $1$
$\lambda_{\ell-2}$ $1$
$\lambda_{\ell-1}$ $\ell-1$
[…]
00 $\ell(\ell+1)/2$
Proof.
We shall present the computations for $i=2$ and we may remark that the case $i=3$ follows very similarly by resorting to the analogous arguments. So, we put $\lambda=\lambda_{\ell-2}$ and $\mu=\lambda_{\ell-1}$. Since $\mu=\lambda-(\alpha_{1}+\alpha_{2}+\cdots+\alpha_{\ell-2})$ we have to consider in the Freudenthal formula the positive roots of the form
$\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j},\qquad 1\leq i\leq j\leq\ell-2.$
Indeed, the formula for $m_{\lambda}(\mu)$ involves the multiplicities of $\mu+(\alpha_{i}+\cdots+\alpha_{j})$ for any $21\leq i\leq j\leq\ell-$ 2. We then observe
$m_{\lambda}(\mu)$ $=2\cdot\frac{\displaystyle{\sum_{1\leq i\leq j\leq\ell-2}m_{\lambda}(\mu+(\alpha_{i}+\cdots+\alpha_{j}))(\mu+(\alpha_{i}+\cdots+\alpha_{j}),\alpha_{i}+\cdots+\alpha_{j})}}{(\lambda,\lambda)-(\mu,\mu)+2(\sum_{k=1}^{\ell-2}\alpha_{k},\delta)}.$
First, by using $(\lambda_{\ell-2},\alpha_{k})=\delta_{(\ell-2)k}$, $(\alpha_{1},\alpha_{1})=1$ and $(\alpha_{k},\alpha_{k+1})=-1$ we deduce that
$(\lambda,\lambda)-(\mu,\mu)$ $=(\lambda_{\ell-2},\lambda_{\ell-2})-\left(\lambda_{\ell-2}-\sum_{k=1}^{\ell-2}\alpha_{k},\lambda_{\ell-2}-\sum_{k=1}^{\ell-2}\alpha_{k}\right)$
$=2\left(\lambda_{\ell-2},\sum_{k=1}^{\ell-2}\alpha_{k}\right)-\left(\sum_{k=1}^{\ell-2}\alpha_{k},\sum_{k=1}^{\ell-2}\alpha_{k}\right)$
$=2\left(\lambda_{\ell-2},\alpha_{\ell-2}\right)-(\alpha_{1},\alpha_{1})-\sum_{k=2}^{\ell-2}(\alpha_{k},\alpha_{k})-2\sum_{k=1}^{\ell-3}(\alpha_{k},\alpha_{k+1})$
$=2-1-2(\ell-3)+2(\ell-3)=1.$
Thus
$(\lambda,\lambda)-(\mu,\mu)+2(\sum_{k=1}^{\ell-2}\alpha_{k},\delta)=1+2(1/2+(\ell-3))=2(\ell-2).$
Now, we claim that the positive roots $\alpha_{1}+\alpha_{2}+\cdots+\alpha_{k}$, with $21\leq k\leq\ell-$ 2, are the only ones that contribute in the computation of $m_{\lambda}(\mu)$. Indeed, we can show that $m_{\lambda}(\mu+(\alpha_{i}+\cdots+\alpha_{j}))=0$, for any $22\leq i\leq j\leq\ell-$ 2. To see that, we directly verify that $m_{\lambda}(\lambda-\alpha_{k})=0$ for any $31\leq k\leq\ell-$ 3 which implies inductively that $m_{\lambda}(\lambda-(\alpha_{1}+\cdots+\alpha_{i-1}))=0$, i.e, we prove the case $j=\ell-2$. Now, by assuming now that $j<\ell-2$ we have $\mu+(\alpha_{i}+\cdots+\alpha_{j})=\lambda-(\alpha_{1}+\cdots+\alpha_{i-1})+(\alpha_{j+1}+\cdots+\alpha_{\ell-2}).$
Then we can note that the kind of positive roots that contribute in the Fruedental formula to compute the foregoing multiplicity are of the form $\alpha_{i^{\prime}}+\cdots+\alpha_{j^{\prime}}$ and $\alpha_{j}$, with $i\leq i^{\prime}\leq j^{\prime}\leq j$, so we may deduce that $m_{\lambda}(\lambda-(\alpha_{1}+\cdots+\alpha_{i-1}+\alpha_{j}))=0$ for any $i+1\leq j\leq\ell-2$. By combining these multiplicities we can conclude that $m_{\lambda}(\lambda-(\alpha_{1}+\cdots+\alpha_{i-1})-(\alpha_{j+1}+\cdots+\alpha_{t}))=0$, for $j+1\leq t\leq\ell-2$. In summary, we have proved that
(5.2) $m_{\lambda}(\mu+(\alpha_{i}+\cdots+\alpha_{j}))=0,\quad 2\leq i\leq j<\ell-2.$
Next by proving
(5.3) $m_{\lambda}(\mu+(\alpha_{1}+\cdots+\alpha_{k}))=1,\quad 1\leq k<\ell-2,$
we can deduce
$m_{\lambda}(\mu)$ $=2\cdot\frac{\displaystyle{\sum_{1\leq i\leq j\leq\ell-2}m_{\lambda}(\mu+(\alpha_{i}+\cdots+\alpha_{j}))(\mu+(\alpha_{i}+\cdots+\alpha_{j}),\alpha_{i}+\cdots+\alpha_{j})}}{(\lambda,\lambda)-(\mu,\mu)+2(\sum_{k=1}^{\ell-2}\alpha_{k},\delta)}$
$=2\cdot\frac{\displaystyle{\sum_{k=0}^{\ell-2}m_{\lambda}(\lambda_{\ell-2}-(\alpha_{k+1}+\cdots+\alpha_{\ell-2}))(\lambda-(\alpha_{k+1}+\cdots+\alpha_{\ell-2}),\alpha_{1}+\cdots+\alpha_{k})}}{2(\ell-2)}$
$=2\cdot\frac{\ell-2}{2(\ell-2)}=1.$
Regarding the proof of the equation (5.3) we proceed by induction. We begin by verifying
$m_{\lambda}(\mu+(\alpha_{1}+\cdots+\alpha_{\ell-4}))=m_{\lambda}(\lambda-(\alpha_{\ell-3}+\alpha_{\ell-2}))=1$
thanks to $m_{\lambda}(\lambda-\alpha_{\ell-2})=1$, $m_{\lambda}(\lambda-\alpha_{\ell-3})=0$ and $m_{\lambda}(\lambda)=1$. Now we assume by induction that
$m_{\lambda}(\mu+(\alpha_{1}+\cdots+\alpha_{k-1}))=m_{\lambda}(\lambda-(\alpha_{k}+\cdots+\alpha_{\ell-2}))=1,\quad 1\leq k<\ell-2,$
and we write $\mu_{k}=\lambda-(\alpha_{k-1}+\alpha_{k}+\cdots+\alpha_{\ell-2})$.
By arguing as above and thanks to (5.2) only the weights of the form $\mu_{k}+(\alpha_{k-1}+\alpha_{k}+\cdots+\alpha_{j})$ will contribute in the Fruedenthal formula of $m_{\lambda}(\lambda-(\alpha_{k-1}+\alpha_{k}+\cdots+\alpha_{\ell-2}))$. So, we observe
$(\mu_{k}+(\alpha_{k-1}+\cdots+\alpha_{j}),\alpha_{k-1}+\cdots+\alpha_{j})=(\lambda-(\alpha_{j+1}+\cdots+\alpha_{\ell-2}),\alpha_{k-1}+\cdots+\alpha_{j})=-(\alpha_{j+1},\alpha_{j})=1.$
Thus,
$m_{\lambda}(\mu_{j})$ $=2\cdot\frac{\displaystyle{\sum_{k=j}^{\ell-2}m_{\lambda}(\mu_{j}+(\alpha_{j}+\cdots+\alpha_{k}))(\mu_{j}+(\alpha_{j}+\cdots+\alpha_{k}),\alpha_{j}+\cdots+\alpha_{k})}}{(\lambda,\lambda)-(\mu_{j},\mu_{j})-2(\sum_{k=j}^{\ell-2}\alpha_{k},\delta)}$
$=2\cdot\frac{\displaystyle{\sum_{k=j}^{\ell-2}m_{\lambda}(\lambda-(\alpha_{k+1}+\cdots+\alpha_{\ell-2}))(\lambda-(\alpha_{k+1}+\cdots+\alpha_{\ell-2}),\alpha_{j}+\cdots+\alpha_{k})}}{(\lambda,\lambda)-(\mu_{j},\mu_{j})-2(\sum_{k=j}^{\ell-2}\alpha_{k},\delta)}$
$=2\cdot\frac{\ell-j-1}{2(\ell-j-1)}=1.$
This completes the computation of $m_{\lambda}(\lambda_{\ell-1})=1$.
Next we shall compute $m_{\lambda}(\lambda_{\ell})$. By observing
$\mu:=\lambda_{\ell}=\lambda_{\ell-2}-\sum_{i=1}^{\ell-2}\alpha_{i}-\sum_{i=1}^{\ell-1}\alpha_{i}$
we can infer that the only terms nonzero in the Freudenthal formula depend on the following positive roots:
$\alpha$ $=\sum_{i=1}^{\ell-2}\alpha_{i}+\sum_{i=1}^{\ell-1}\alpha_{i},$
[…]
$\alpha_{k,j}$ $=2(\alpha_{1}+\cdots+\alpha_{k})+(\alpha_{k+1}+\cdots+\alpha_{j}),\quad 1\leq k\leq\ell-3\;\text{and }\ell-2\leq j\leq\ell-1.$
At this point, we need to remark that $m_{\lambda}(\mu+2(\alpha_{1}+\cdots+\alpha_{k}))(\mu+2(\alpha_{1}+\cdots+\alpha_{k}),(\alpha_{1}+\cdots+\alpha_{k}))$ also contributes in the Freudenthal formula of $m_{\lambda}(\mu)$, when $31\leq k\leq\ell-$ 3. According to the terms associated to the above positive roots in the foregoing formula, we first compute that
$m_{\lambda}\left(\mu+\sum_{i=1}^{\ell-2}\alpha_{i}+\sum_{i=1}^{\ell-1}\alpha_{i}\right)\cdot\left(\mu+\sum_{i=1}^{\ell-2}\alpha_{i}+\sum_{i=1}^{\ell-1}\alpha_{i},\sum_{i=1}^{\ell-2}\alpha_{i}+\sum_{i=1}^{\ell-1}\alpha_{i}\right)=m_{\lambda}(\lambda)\left(\lambda,\sum_{i=1}^{\ell-2}\alpha_{i}+\sum_{i=1}^{\ell-1}\alpha_{i}\right)=1\cdot 2.$
Supposing that $\alpha_{k}=\alpha_{1}+\cdots+\alpha_{k}$, where $11\leq k\leq\ell-$ 1, we directly verify that $(\mu+\alpha_{k},\alpha_{k})=1$. Now we shall prove that $m_{\lambda}(\mu+\alpha_{k})=1$. Let us start by noticing that the multiplicities for $k=\ell-i,\;i=1,2$ were computed in the above case, thus we have to compute $m_{\lambda}(\mu+\alpha_{k})$ for $k\leq\ell-3$. For the general case, the Freudenthal formula involves computing many terms of the form $m_{\lambda}(\mu_{k}+\alpha))(\mu_{k}+\alpha,\alpha)$ for appropriate positive roots $\alpha$ we start by setting $\mu_{k}=\mu+\alpha_{k}$. Each of them implies in its turn to compute many other multiplicities. Nevertheless, we introduce a case sufficiently illustrative of these computations, say $m_{\lambda}(\mu_{\ell-3})$. Firstly, we see that
$\mu_{\ell-3}=\lambda_{\ell-2}-(\alpha_{1}+\cdots+\alpha_{\ell-3}+2\alpha_{\ell-2}+\alpha_{\ell-1}).$
Immediately, we calculate
$(\lambda,\lambda)-(\mu_{\ell-3},\mu_{\ell-3})+2(\alpha_{\ell-1},\delta)=3-3+2\ell=2\ell.$
Now by appealing to similar arguments and strategies as in the computing of $m_{\lambda_{\ell-2}}(\lambda_{\ell-1})$ we may deduce the only positive roots contributing to the formula to compute $m_{\lambda}(\mu_{\ell-3})$ are
[…]
Here we also argue as in $m_{\lambda}(\mu_{\ell-1})$ we may obtain
$m_{\lambda}(\mu+\alpha_{1}+\cdots+\alpha_{k})(\mu+\alpha_{1}+\cdots+\alpha_{\ell-1},\alpha_{1}+\cdots+\alpha_{k})$ $=1\cdot 1,\quad 1\leq k\leq\ell-1,\;\;k\neq\ell-3,$
$m_{\lambda}(\mu+\alpha_{1}+\cdots+\alpha_{\ell-3})(\mu+\alpha_{1}+\cdots+\alpha_{\ell-1},\alpha_{1}+\cdots+\alpha_{\ell-3})$ $=1\cdot 2.$
Therefore, we get
$m_{\lambda}(\mu)=\frac{2[(\ell-2)\cdot 1+1\cdot 2]}{2\ell}=1.$
Regarding the term $m_{\lambda}(\mu+2\alpha_{k})(\mu+2\alpha_{k},\alpha_{k})$, for $31\leq k\leq\ell-$ 3, we can proceed similarly as above to deduce each multiplicity $m_{\lambda}(\mu+2\alpha_{k})$. At this point, routine computations allow to calculate the remaining components to obtain the multiplicities.
Finally, suppose that $\alpha_{k,j}=2(\alpha_{1}+\cdots+\alpha_{k})+(\alpha_{k+1}+\cdots+\alpha_{j})$, where $31\leq k\leq\ell-$ 3 and $\ell-2\leq j\leq\ell-1$. We compute $(\mu+\alpha_{k,j},\alpha_{k,j})$ as follows:
  • For $j=\ell-1$ we have
    $(\mu+\alpha_{k,\ell-1},\alpha_{k,\ell-1})$ $=(\lambda_{\ell-2}-(\alpha_{k+1}+\cdots+\alpha_{\ell-2}),2(\alpha_{1}+\cdots+\alpha_{k})+(\alpha_{k+1}+\cdots+\alpha_{\ell-1}))$
    $=(\lambda_{\ell-2},\alpha_{\ell-2})-2(\alpha_{k+1},\alpha_{k})-(\alpha_{k+1},\alpha_{k+1})-(\alpha_{\ell-2},\alpha_{\ell-1})$
    $=2.$
    Likewise we verify that $(\mu+\alpha_{k,\ell-2},\alpha_{k,\ell-2})=2$.
  • When $k\leq\ell-3$ we get
    $(\mu+\alpha_{k,\ell-1},\alpha_{k,\ell-1})$ $=(\lambda_{\ell-2}-(\alpha_{k+1}+\cdots+\alpha_{\ell-2}),2(\alpha_{1}+\cdots+\alpha_{k})+(\alpha_{k+1}+\cdots+\alpha_{\ell-1}))$
    $=(\lambda_{\ell-2},\alpha_{\ell-2})-2(\alpha_{k+1},\alpha_{k})-(\alpha_{k+1},\alpha_{k+1})-(\alpha_{\ell-2},\alpha_{\ell-1})$
    $=2,$
    and identically we get that $(\mu+\alpha_{k,\ell-2},\alpha_{k,\ell-2})=2$.
On the other hand, we may resort as above to determine the only positive roots $\alpha$ involving in the formula of $m_{\lambda}(\mu+\alpha_{k,j})$ we recursively deduce that $m_{\lambda}(\lambda_{\ell-2}-(\alpha_{k+1}+\cdots+\alpha_{\ell-2}))=1$ as in the previous case. Since $31\leq k\leq\ell-$ 3 and $\ell-2\leq j\leq\ell-1$ it follows that the number of positive roots $\alpha_{k,j}$ is
$1\cdot 2+2\cdot 2+\cdots+2\cdot(\ell-3)=2\cdot(\ell-2)(\ell-3)/2=(\ell-2)(\ell-3).$
In summary, the computations from the cases above provide respectively the following summands to the formula of $m_{\lambda}(\mu)$
12,(1)11,(3)12,(2)12, $1\cdot 2,\;(\ell-1)\cdot 1\cdot 1,\;(\ell-3)\cdot 1\cdot 2,\;(\ell-2)\cdot 1\cdot 2,\;1\cdot 2(\ell-2)(\ell-3)$
Then we can compute
$m_{\lambda}(\mu)$ $=\frac{2[2\cdot 1+(\ell-1)\cdot 1+2(\ell-2)+2(\ell-3)+2(\ell-2)(\ell-3)]}{2(2\ell-3)}$
$=\frac{2(\ell-1)(2\ell-3)}{2(2\ell-3)}$
$=\ell-1.$
To obtain the multiplicity $m_{\lambda}(0)$, we observe that
$\mu=0=\lambda-\left(\sum_{i=2}^{\ell-2}\alpha_{\ell-i}+\sum_{i=1}^{\ell-1}\alpha_{\ell-i}+\sum_{i=0}^{\ell}\alpha_{\ell-i}\right).$
This writing allows to apply the same strategy as above to determine all terms involved in the formula of $m_{\lambda}(0)$ and thereby to conclude the desired multiplicity.
The case $\lambda=4\lambda_{\ell}$ is only longer, but completely similar than the ones considered above. The natural first step is to compute the multiplicity of $\lambda_{\ell-1}+2\lambda_{\ell}$. ∎
Lemma 5.5.
For all $\ell\geq 12$, the multiplicities of all dominant weights $\mu\preceq\lambda_{\ell-1}+\lambda_{\ell}$ are given in the following table:
highest weight dominant weight $\mu\preceq\lambda$ multiplicity of $\mu$
$\lambda=\lambda_{\ell-1}+\lambda_{\ell}$ $\lambda_{\ell-1}+\lambda_{\ell}$ $1$
$\lambda_{\ell-2}$ $2$
$2\lambda_{\ell}$ $1$
$\lambda_{\ell-1}$ $2$
$\lambda_{\ell}$ $12\ell-$ 1
00 $12\ell-$ 1
Proof.
We shall start by computing the multiplicity of $\lambda_{\ell-2}=\lambda-(\alpha_{\ell-1}+\alpha_{\ell})$. To do this, we can verify as above that
$m_{\lambda}(\lambda-\alpha_{\ell})=2\cdot\frac{m_{\lambda}((\lambda-\alpha_{\ell})+\alpha_{\ell})((\lambda-\alpha_{\ell})+\alpha_{\ell},\alpha_{\ell})}{(\lambda,\lambda)-(\mu_{\ell},\mu_{\ell})+2(\alpha_{\ell},\delta)}=2\cdot\frac{1\cdot 1}{5-5+2\cdot 1}=1.$
and similarly we check that $m_{\lambda}(\lambda-\alpha_{\ell-1})=1$. Thus, we may deduce that the multiplicity $m_{\lambda}(\lambda_{\ell-2})=2$ by computing as follows
$m_{\lambda}(\lambda_{\ell-2})$ $=2\cdot\frac{m_{\lambda}(\mu_{\ell-1}+\alpha_{\ell-1})(\mu_{\ell-1}+\alpha_{\ell-1},\alpha_{\ell-1})+m_{\lambda}(\mu_{\ell-1}+\alpha_{\ell})(\mu_{\ell-1}+\alpha_{\ell},\alpha_{\ell})+m_{\lambda}(\lambda)(\lambda,\alpha_{\ell-1}+\alpha_{\ell})}{(\lambda,\lambda)-(\lambda_{\ell-2},\lambda_{\ell-2})+2(\alpha_{\ell-1}+\alpha_{\ell},\delta)}$
$=2\cdot\frac{1\cdot 2+1\cdot 2+1\cdot 2}{5-3+4}=2.$
In order to compute $m_{\lambda}(2\lambda_{\ell})$ we see that
$\mu:=2\lambda_{\ell}=\lambda-\sum_{k=1}^{\ell-1}\alpha_{k}.$
A direct computation shows
$(\lambda,\lambda)-(\mu,\mu)+2(\sum_{k=1}^{\ell-1}\alpha_{k},\delta)=5-4+2(1/2+(\ell-2))=2(\ell-1).$
By arguing as in proof of Lemma 5.3, we may infer that the only required multiplicities to compute $m_{\lambda}(2\lambda_{\ell})$ are $m_{\lambda}(\mu+(\alpha_{1}+\cdots+\alpha_{k}))$, for $21\leq k\leq\ell-$ 2, and $m_{\lambda}(\lambda)$ itself. Inductively, as Lemma 5.4, we may check that $m_{\lambda}(\mu+(\alpha_{1}+\cdots+\alpha_{k}))=1$ for any $21\leq k\leq\ell-$ 2. In addition, we can also verify in this case that
$(\mu+(\alpha_{1}+\cdots+\alpha_{k}),\alpha_{1}+\cdots+\alpha_{k})$ $=(\lambda-(\alpha_{k+1}+\cdots+\alpha_{\ell-1}),\alpha_{1}+\cdots+\alpha_{k})=-(\alpha_{k+1},\alpha_{k})=1.$
for $k\leq\ell-2$. Thus, we deduce
$m_{\lambda}(\mu)$ $=2\cdot\frac{\displaystyle{\sum_{k=1}^{\ell-1}m_{\lambda}(\mu+(\alpha_{1}+\cdots+\alpha_{k}))(\mu+(\alpha_{1}+\cdots+\alpha_{k}),\alpha_{1}+\cdots+\alpha_{k})}}{(\lambda,\lambda)-(\mu,\mu)-2(\sum_{k=1}^{\ell-1}\alpha_{k},\delta)}$
$=2\cdot\frac{\displaystyle{\sum_{k=1}^{\ell-1}m_{\lambda}(\lambda-(\alpha_{k+1}+\cdots+\alpha_{\ell-2}))(\lambda-(\alpha_{k+1}+\cdots+\alpha_{\ell-2}),\alpha_{1}+\cdots+\alpha_{k})}}{2(\ell-1)}$
$=2\cdot\frac{\ell-1}{2(\ell-1)}=1.$
We now start the computations of $m_{\lambda}(\lambda_{\ell-1})$. We first note that
$\mu:=\lambda_{\ell-1}=\lambda-\sum_{k=1}^{\ell}\alpha_{k}.$
We have
$(\lambda,\lambda)-(\mu,\mu)+2(\sum_{k=1}^{\ell}\alpha_{k},\delta)=5-2+2(1/2+(\ell-1))=2(\ell+1).$
In addition we already know that $m_{\lambda}(\lambda-(\alpha_{\ell-1}+\alpha_{\ell}))=2$ and $m_{\lambda}(\lambda-\alpha_{\ell})=1$ by calculations above. Thus, we see that $m_{\lambda}(\lambda)(\lambda,\alpha_{1}+\cdots+\alpha_{\ell})=1\cdot 2$. Indeed, we may argue as in Lemma 5.3 to deduce
$m_{\lambda}(\lambda-(\alpha_{k+1}+\cdots+\alpha_{\ell}))(\lambda-(\alpha_{k+1}+\cdots+\alpha_{\ell}),\alpha_{1}+\cdots+\alpha_{k})=2\cdot 1,$
for any $21\leq k\leq\ell-$ 2, and $m_{\lambda}(\lambda)(\lambda,\alpha_{1}+\cdots+\alpha_{\ell})=1\cdot 2$. In the same way, we can verify
$m_{\lambda}(\lambda-(\alpha_{1}+\cdots+\alpha_{\ell-1}))(\lambda-(\alpha_{1}+\cdots+\alpha_{\ell-1}),\alpha_{\ell})=1\cdot 2.$
Therefore, by noting that the foregoing terms are the only nonzero summands in the Fruedenthal formula as in Lemma 5.4 we conclude that $m_{\lambda}(\lambda_{\ell-1})=2$.
While the multiplicity $m_{\lambda}(\lambda_{\ell})$ may be computed by means of the arguments used in Lemma 5.4 observing that
$\mu:=\lambda_{\ell}=\lambda-\sum_{k=1}^{\ell-1}\alpha_{k}-\sum_{k=1}^{\ell}\alpha_{k}=\lambda-2\sum_{k=1}^{\ell-1}\alpha_{k}-\alpha_{\ell}$
and
$(\lambda,\lambda)-(\mu,\mu)+2(2\sum_{k=1}^{\ell-1}\alpha_{k}+\alpha_{\ell},\delta)=5-1+4(\ell-1)=4\ell.$
Idem for $m_{\lambda}(0)$ taking in mind
$\mu:=0=\lambda-\sum_{k=1}^{\ell-1}\alpha_{k}-2\sum_{k=1}^{\ell}\alpha_{k}=3(2\ell-1).$
and
$(\lambda,\lambda)-(\mu,\mu)+2(\sum_{k=1}^{\ell-1}\alpha_{k}+2\sum_{k=1}^{\ell}\alpha_{k},\delta)=5-0+3(2\ell-1))=2(3\ell+1).$
Thus we can prove.
Proposition 5.6.
Let $G$ be as above of type $B_{\ell}$ with $\ell\geq 12$. Then, all weights in the following table are admissible.
$\lambda$ dimension of $V(\lambda)$
00 $1$
$\lambda_{\ell}$ $12\ell+$ 1
$2\lambda_{\ell}$ $\ell(2\ell+3)-1$
$3\lambda_{\ell}$ $\ell(2\ell+1)(2\ell+5)/3$
$4\lambda_{\ell}$ $\ell(\ell+1)(2\ell+1)(2\ell+7)/6$
$\lambda_{\ell-1}$ $2\ell^{2}+\ell$
$\lambda_{\ell-2}$ $(2\ell-1)2\ell(2\ell+1)/6$
$\lambda_{\ell-3}$ $(2\ell-2)(2\ell-1)2\ell(2\ell+1)/24$
$\lambda_{\ell}+\lambda_{\ell-1}$ $(2\ell+3)(2\ell+1)(2\ell-1)/3$
In addition, when $12\leq\ell\leq 15$ the table has to include the weight $\lambda_{1}$ whose associated irreducible representation $V(\lambda_{1})$ has dimension $2^{\ell}$.
Proof.
Suppose that $\ell\geq 12$. The dimensions of the weight spaces associated to the highest weights 00, $\lambda_{\ell}$, and $2\lambda_{\ell}$, are computed in Table 2 of [11]. We next shall compute the dimension of $V(3\lambda_{\ell})$, the dimension of $V(4\lambda_{\ell})$ follows by means of the same computations. To this end and according to Proposition 2.1 we have to compute $l(\mu)$, for each dominant weight $\mu\preceq 3\lambda_{\ell}$ given in Lemma 5.3. We first observe that the stabilizers of the weights $\mu_{i}=i\lambda_{\ell}$, for $i=1,2,3$, are reflection groups of type $B_{\ell-1}$, and hence $l(\mu_{i})=2\ell$. On the other hand, we also see that the stabilizers of $\lambda_{\ell-1}+\lambda_{\ell}$, $\lambda_{\ell-1}$ and $\lambda_{\ell-2}$ are reflection groups of type $B_{\ell-2}$, $B_{\ell-2}\times A_{1}$ and $B_{\ell-3}\times A_{2}$, so the length of its orbits are $4\ell(\ell-1)$, $2\ell(\ell-1)$, $34\ell(\ell-1)(\ell-2)/$ 3 respectively and clearly $l(0)=1$. Then, by resorting to the multiplicities given in Lemma 5.3 and Proposition 2.1 we get
$\dim V(3\lambda_{\ell})$ $=2\ell+4\ell(\ell-1)+4\ell(\ell-1)(\ell-2)/3+2\ell+2\ell(\ell-1)+2\ell^{2}+\ell$
$=\ell(2\ell+1)(2\ell+5)/3.$
Now we compute the dimension of $V(\lambda_{\ell-3})$ since the case of $V(\lambda_{\ell-2})$ follows in the same way, even shorter, and $V(\lambda_{\ell-1})$ was computed in [11]. Certainly $\dim V(\lambda_{\ell-3})$ may be calculated as above, but let us give an alternative way to deduced it. We first observe that
$\dim V(\lambda_{\ell-3})\geq m_{\lambda_{\ell-3}}(\lambda_{\ell-3})l(\lambda_{\ell-3})+m_{\lambda_{\ell-3}}(\lambda_{\ell-2})l(\lambda_{\ell-2})+m_{\lambda_{\ell-3}}(\lambda_{\ell-1})l(\lambda_{\ell-1})+m_{\lambda_{\ell-3}}(\lambda_{\ell})l(\lambda_{\ell})+m_{\lambda_{\ell-3}}(0)l(0).$
As above we see that the stabilizer of $\lambda_{\ell-i}$ is a reflection group of type $B_{\ell-i-1}\times A_{i}$, for $i=1,2,3$ and of type $B_{\ell-2}$ for $\lambda_{\ell}$. Then, $l(\mu)$ is 2(1)(2)(3)/3, 4(1)(2)/3, 2(1), 2, $12\ell(\ell-1)(\ell-2)(\ell-3)/3,\>4\ell(\ell-1)(\ell-2)/3,\;2\ell(\ell-1),\;2\ell,\;$ 1 for $\mu=\lambda_{\ell-3},\;\mu=\lambda_{\ell-2},\;\mu=\lambda_{\ell-1},\;\mu=\lambda_{\ell},\;\mu=0$, respectively. Thus, by using the multiplicities given by Lemma 5.3 we see that
$\dim V(\lambda_{\ell-3})$ $\geq 2\ell(\ell-1)(\ell-2)(\ell-3)/3+4\ell(\ell-1)(\ell-2)/3+2\ell(\ell-1)(\ell-2)+2\ell(\ell-1)+\ell(\ell-1)/2$
$=(2\ell-2)(2\ell-1)2\ell(2\ell+1)$
$=\dim\Lambda^{4}(V).$
By combining the last inequality with $V(\lambda_{\ell-3})\subseteq\Lambda^{4}(V)$ we conclude that $V(\lambda_{\ell-3})$ is equal to $\Lambda^{4}(V)$.
On the other hand, we can deduce the dimension of $V(\lambda_{\ell-1}+\lambda_{\ell})$ by appealing Lemma 5.5 and arguing as in $V(3\lambda_{\ell})$.
Finally, for $\lambda=\lambda_{1}$ and $12\leq\ell\leq 15$ we infer that there not exist $\mu\preceq\lambda$ except $\lambda$ itself. In addition, we directly verify that $\rm{Orb}_{\mathcal{W}}(\lambda)=2^{\ell}$. This then says that $\dim V(\lambda)=2^{\ell}<\ell^{4}$ for $\ell$ in the mentioned rank, i.e., $\lambda$ is admissible. ∎
Remark 5.7.
Let $G$ be as above of type $B_{\ell}$ and $V$ be the irreducible $G$-module of dimension $12\ell+$ 1. Then, the anti-symmetric tensor product $\Lambda^{i}(V)$ is irreducible, for $i=1,2,3,4$.
In order to prove that the weights in the previous proposition are the only admissible weights of type $B_{\ell}$, we will prove the following combinatorial lemma which will be very useful below. This result is perhaps well-known to the experts, but we give a proof for lack of a reference.
Lemma 5.8.
The coefficients of
$(2+x)^{n}=\sum_{k=0}^{n}\binom{n}{k}2^{n-k}x^{k}$
for $n\geq 12$ are strictly increasing until $k=\lfloor\frac{n}{3}\rfloor$ and then they are strictly decreasing.
Proof.
Firstly we put $t_{k}:=2^{n-k}\binom{n}{k}$ for each integer $0\leq k\leq n$. Then
$t_{k}:=2^{n-k}\binom{n}{k}=2\cdot\frac{2^{n-(k+1)}n(n-1)\cdots(n-(k+1))}{k!}.$
We begin by observing $t_{0}=2^{n-1}\cdot 2<2^{n-1}\cdot n=t_{1}$ because $2<n$. We also note that
$t_{1}=2^{n-1}n=2^{n-2}n\cdot 2<2^{n-2}n\cdot\frac{n-1}{2}=t_{2},$
since $22<(n-1)/$ 2. Thus, $t_{0}<t_{1}<t_{2}$. In general, if $k<\lfloor\frac{n}{3}\rfloor$ we may infer that $2k<(n-(k+1))$. It works for $n\geq 12$ and proceed as above to show $t_{k}<t_{k+1}$. We see that the worst case is when $k=\lfloor\frac{n}{3}\rfloor-1$. In such case, the inequality follows from
$n-\left(\left\lfloor\frac{n}{3}\right\rfloor-1\right)\geq n-\frac{n}{3}+1=2\frac{n}{3}+1\geq 2\left\lfloor\frac{n}{3}\right\rfloor.$
We then deduce
$t_{\lfloor\frac{n}{3}\rfloor-1}$ $=2^{n-\lfloor\frac{n}{3}\rfloor+1}\binom{n}{\lfloor\frac{n}{3}\rfloor-1}$
$=2^{n-\lfloor\frac{n}{3}\rfloor+1}\frac{n!}{(\lfloor\frac{n}{3}\rfloor-1)!(n-\lfloor\frac{n}{3}\rfloor+1)!}$
$=2^{n-\lfloor\frac{n}{3}\rfloor+1}\frac{n!}{(\lfloor\frac{n}{3}\rfloor-1)!(n-\lfloor\frac{n}{3}\rfloor)!(n-\lfloor\frac{n}{3}\rfloor+1)}$
$=2^{n-\lfloor\frac{n}{3}\rfloor}\frac{n!}{(\lfloor\frac{n}{3}\rfloor-1)!(n-\lfloor\frac{n}{3}\rfloor)!}\cdot\frac{2}{(n-\lfloor\frac{n}{3}\rfloor+1)}$
$<2^{n-\lfloor\frac{n}{3}\rfloor}\frac{n!}{(\lfloor\frac{n}{3}\rfloor-1)!(n-\lfloor\frac{n}{3}\rfloor)!}\cdot\frac{1}{\lfloor\frac{n}{3}\rfloor}$
$=t_{\lfloor\frac{n}{3}\rfloor},$
thanks to the inequality $2/(n-\lfloor\frac{n}{3}\rfloor+1)<1/\lfloor\frac{n}{3}\rfloor$. This proves that the coefficients of $(2+x)^{n}$ are strictly increasing until $k=\lfloor\frac{n}{3}\rfloor$.
On the other hand, by similar arguments, when $k>\lfloor\frac{n}{3}\rfloor$ we have that $2\geq\frac{n-(k-1)}{k}$. Then, by observing,
$2^{n-(k+1)}\frac{n(n-1)\cdots(n-(k-2))}{(k-1)!}\cdot 2\geq 2^{n-k}\frac{n(n-1)\cdots(n-(k-2))}{(k-1)!}\cdot\frac{(n-(k-1))}{k}$
we show that the coefficients of $(2+x)^{n}$ are strictly decreasing when $k>\lfloor\frac{n}{3}\rfloor$. This finishes the proof. ∎
Now we are ready to state and give a proof of our main result of this section.
Theorem 5.9.
Let $G$ be a finite group of Lie type $B_{\ell}$ over $\mathbb{\overline{F}}_{p}$. If $\ell\geq 12$ then the list giving in Proposition 5.6 exhausts the admissible weights.
Proof.
We shall proceed case-by-case by arguing as in Theorem 4.1. Start by setting
$\lambda=a_{1}\lambda_{1}+a_{2}\lambda_{2}+\cdots+a_{\ell}\lambda_{\ell}.$
Nullity of coefficients -- $\lambda$ is admissible implies $a_{i}=0$ for $41\leq i\leq\ell-$ 4.
Suppose first that $a_{i}\neq 0$, for some $42\leq i\leq\ell-$ 4. Then, the stabilizer of $\lambda$ is contained in a reflection group of type $B_{i-1}\times A_{\ell-i}$ it is worth mentioning that we are using the convention $B_{1}=A_{1}$ as in [6]. This implies
$\dim V(\lambda)\geq t_{i}:=\frac{2^{\ell}\ell!}{2^{i-1}(i-1)!(\ell-i+1)!}$ $=2^{\ell-(i-1)}\binom{\ell}{i-1}.$
By appealing to Lemma 5.8 we have that the $(i-1)$th-coefficient $t_{i}$ of $(2+x)^{\ell}$ are strictly increasing for $12\leq i\leq\lfloor\frac{\ell}{3}\rfloor+$ 1. Thus, for $\ell\geq 16$ we use the fact that $2^{\ell}\geq\ell^{4}$ to deduce
$\dim(V(\lambda))\geq t_{i}>t_{0}=2^{\ell}\geq\ell^{4}.$
Now, by numerical inspection we also may verify for each $12\leq\ell\leq 15$ that
$\dim V(\lambda)\geq$ $1\displaystyle 2^{\ell-(i-1)}\binom{\ell}{i-1}>\ell^{4},\quad\quad\text{ for }2\leq i\leq\lfloor\frac{\ell}{3}\rfloor+$ 1 .
On the other hand, we will demonstrate that $\dim V(\lambda)>\ell^{4}$ by using the sequence $(t_{i})$ is decreasing for $i>\lfloor\frac{\ell}{3}\rfloor+1$. Indeed, we shall show that $t_{\ell-4}>\ell^{4}$ for $\ell\geq 12$ by observing
$t_{\ell-4}>\ell^{4}$ $\iff 2^{5}\frac{\ell(\ell-1)(\ell-2)(\ell-3)(\ell-4)}{5!}>\ell^{4}$
$\iff 2^{5}\frac{(\ell-1)(\ell-2)(\ell-3)(\ell-4)}{5!}>\ell^{3}$
$\iff 2^{5}\frac{\ell^{3}(1-1/\ell)(1-2/\ell)(1-3/\ell)(\ell-4)}{2\cdot 3\cdot 4\cdot 5}>\ell^{3}$
$\iff 2^{2}\frac{(1-1/\ell)(1-2/\ell)(1-3/\ell)(\ell-4)}{3\cdot 5}>1.$
Since $\ell\geq 12$ we can infer that $121-j/\ell\geq(12-j)/$ 12 for $1\leq j\leq 3$, and $(\ell-4)/15\geq 8/15$. By combining these facts we can infer
$2^{2}\frac{(1-1/\ell)(1-2/\ell)(1-3/\ell)(\ell-4)}{3\cdot 5}$ $=2^{2}(1-1/\ell)(1-2/\ell)(1-3/\ell)(\ell-4)/15$
$\geq 2^{2}\cdot\frac{11}{12}\cdot\frac{10}{12}\cdot\frac{9}{12}\cdot\frac{8}{15}>1.$
With the foregoing equivalence and inequality above, we can conclude that $t_{\ell-4}>\ell^{4}$ for $\ell\geq 12$. This implies that $t_{i}\geq t_{\ell-4}>\ell^{4}$, for $i\geq\lfloor\frac{\ell}{3}\rfloor+1$, when $\ell\geq 12$, by appealing to the sequence $(t_{i})$ is decreasing in this rank. We then conclude
$\dim(V(\lambda))\geq t_{i}\geq t_{\ell-4}>\ell^{4},$
To deal the case $i=1$ we proceed as follows. If $a_{1}\geq 2$ we deduce that $\lambda^{\prime}:=(\lambda-\alpha_{\ell-1})\prec\lambda$ can be written as nonnegative linear combination of fundamental weights contains the term $\lambda_{2}$. Resorting as above we have that
$\dim(V(\lambda))\geq\frac{|\mathcal{W}|}{|\rm{Stab}_{\mathcal{W}}(\lambda^{\prime})|}>\ell^{4},$
i.e., $\lambda$ is not admissible when $a_{1}\geq 2$.
By appealing to the same arguments, $a_{1}=1$ implies $a_{i}=0$ for any $2\leq i\leq\ell$. Hence we reduce to
$\lambda=\lambda_{1}\quad\quad\text{ or }\quad\quad\lambda=a_{\ell-3}\lambda_{\ell-3}+a_{\ell-2}\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}.$
By Proposition 5.6 we have that $\lambda=\lambda_{1}$ is admissible when $12\leq\ell\leq 15$ and not otherwise. Next we consider $\lambda=a_{\ell-3}\lambda_{\ell-3}+a_{\ell-2}\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$. If $a_{\ell-3}\geq 2$, we produce a dominant weight $\lambda^{\prime}:=\lambda-\alpha_{\ell-3}$, whose writing as nonnegative linear combination of fundamental weights contains the term $\lambda_{\ell-4}$, but it cannot be possible due to the arguments above. Thus, we reduce the analysis to
$\lambda=\lambda_{\ell-3}+a_{\ell-2}\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}\;\text{ or }\;\lambda=a_{\ell-2}\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}.$
For the former case, i.e., assuming $a_{\ell-3}=1$, we claim that $a_{\ell-j}=0$ for $j=0,1,2$. In fact, if some of these coefficient were nonzero then the stabilizer $\rm{Stab}_{\mathcal{W}}(\lambda)$ of $\lambda$ were contained in a reflection group of type $B_{\ell-4}\times A_{1}$, $B_{\ell-4}\times A_{1}\times A_{1}$ or $B_{\ell-4}\times A_{2}$. In any of those cases we can verify that
$\dim V(\lambda)\geq\frac{|\mathcal{W}|}{|\rm{Stab}_{\mathcal{W}}(\lambda)|}>\ell^{4}.$
by arguing as cases above.
On the other hand, when $\lambda=a_{\ell-2}\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$ we can prove that $a_{\ell-2}<2$. Otherwise, we observe that $\lambda_{\ell-3}$ and $\lambda_{\ell-1}$ appear in the writing of $\lambda-\alpha_{\ell-2}$, but this case was already covered previously. Thus, we get
$\lambda=\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}\quad\text{ or }\quad\lambda=a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}.$
In both cases we may deduce that $\lambda$ is not admissible if $a_{\ell-i}>1$ with $i=0$ or $1$ by arguing as above. For instance, if $\lambda=\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$. with $a_{\ell-1}\geq 2$ we can observe that the term $2\lambda_{\ell-2}$ appears in $\lambda^{\prime}:=\lambda-\alpha_{\ell-1}$, but such a weight $\lambda^{\prime}$ is not admissible by the previous case. It follows that $\lambda=\lambda_{\ell-2}+a_{\ell}\lambda_{\ell}$ or $\lambda=\lambda_{\ell-2}+\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$. Moreover, we may proceed very similarly for $a_{\ell}$ which implies to analyze the dimensions of the representations associated to the weights
$\lambda_{\ell-2}+\lambda_{\ell-1},\;\lambda_{\ell-2}+\lambda_{\ell}\;\text{ and }\lambda_{\ell-2}+\lambda_{\ell-1}+\lambda_{\ell}$
For $\lambda:=\lambda_{\ell-2}+\lambda_{\ell-1}$, by using the identity,
$\alpha_{\ell-2}+\alpha_{\ell-1}=-\lambda_{\ell-3}+\lambda_{\ell-2}+\lambda_{\ell-1}-\lambda_{\ell}$
we deduce that $\mu:=\lambda_{\ell-3}+\lambda_{\ell}\prec\lambda_{\ell-2}+\lambda_{\ell-1}$. The same calculations as above show that the length of orbit of $\mu$ is greater to $\ell^{4}$, and hence $\dim V(\lambda)>\ell^{4}$, i.e., $\lambda$ is not admissible.
On the other hand, $\lambda:=\lambda_{\ell-2}+\lambda_{\ell}$ is not admissible by Example 3.4. While for $\lambda:=\lambda_{\ell-2}+\lambda_{\ell-1}+\lambda_{\ell}$ one notices that
$(\lambda_{\ell-2}+\lambda_{\ell-1}+\lambda_{\ell})-(\lambda_{\ell-3}+2\lambda_{\ell})=-\lambda_{\ell-3}+\lambda_{\ell-2}+\lambda_{\ell-1}-\lambda_{\ell}=\alpha_{\ell-2}+\alpha_{\ell-1}$
which says $\lambda_{\ell-3}+2\lambda_{\ell}\prec\lambda$. Thus, we have
$\dim V(\lambda)$ $\geq l(\lambda_{\ell-3}+2\lambda_{\ell})$
$=2^{3}\ell(\ell-1)(\ell-2)(\ell-3)/3$
$=\ell^{4}\cdot 2^{3}(1-1/\ell)(1-2/\ell)(1-3/\ell)/3$
$>\ell^{4},$
by observing $12^{3}(1-1/\ell)(1-2/\ell)(1-3/\ell)/3>$ 1 for any $\ell\geq 12$. This implies that $\lambda=\lambda_{\ell-2}+\lambda_{\ell-1}+\lambda_{\ell}$ is not admissible either.
Finally, we consider the case $\lambda=a_{\ell}\lambda_{\ell}$. By appealing to Lemma 5.6, the weights of the form $\lambda=a_{\ell}\lambda_{\ell}$ are admissible for $a_{\ell}=1,2,3,4$. To complete this case we observe that none $\lambda$ with $a_{\ell}>4$ provides a new admissible weight. To see that, it suffices to prove
$\lambda_{\ell-4}\prec\lambda_{\ell-3}+\lambda_{\ell}\prec 5\lambda_{\ell}\preceq\lambda.$
forcing then $\dim V(\lambda)\geq l(\lambda_{\ell-4})>\ell^{4}$.
Therefore, the Table in Proposition 5.6 exhausts the admissible weights which conclude the proof. ∎
Finally, we will deal with the rest of the classic groups Lie type in the next subsection.
5.2. Type $C_{\ell}$ and $D_{\ell}$
Theorem 5.10.
Let $G$ be a group of type $C_{\ell}$ or $D_{\ell}$ over $\mathbb{\overline{F}}_{p}$. If $\ell\geq 12$ then the lists giving in the table here below exhaust the admissible weights in each case.
type $\lambda$ dimension
$C_{\ell}$
0 1
$\lambda_{\ell}$ $2\ell$
$2\lambda_{\ell}$ $(2\ell+1)(2\ell)/2$
$3\lambda_{\ell}$ $(2\ell+2)(2\ell+1)(2\ell)/6$
$4\lambda_{\ell}$ $(2\ell+3)(2\ell+2)(2\ell+1)(2\ell)/24$
$\lambda_{\ell-1}$ $12\ell^{2}-\ell-$ 1
$\lambda_{\ell-2}$ $2\ell(2\ell-1)(2\ell-2)/6-2\ell$
$\lambda_{\ell-3}$ $((2\ell-3)(2\ell-2)(2\ell-1)2\ell/24)-((2\ell+1)(2\ell)/2)$
$\lambda_{\ell}+\lambda_{\ell-1}$ $(2\ell+2)(2\ell)(2\ell-2)/3$
$D_{\ell}$
0 1
$\lambda_{\ell}$ $2\ell$
$2\lambda_{\ell}$ $(2\ell+1)(2\ell)/2-1$
$3\lambda_{\ell}$ $(2\ell+2)(2\ell+1)(2\ell)/6-2\ell$
$4\lambda_{\ell}$ $((2\ell+3)(2\ell+2)(2\ell+1)(2\ell)/24)-((2\ell+1)(2\ell)/2)$
$\lambda_{\ell-1}$ $2\ell^{2}-\ell$
$\lambda_{\ell-2}$ $62\ell(2\ell-1)(2\ell-2)/$ 6
$\lambda_{\ell-3}$ $(2\ell-3)(2\ell-2)(2\ell-1)2\ell/24$
$\lambda_{\ell}+\lambda_{\ell-1}$ $(2\ell+2)(2\ell)(2\ell-2)/3$
In addition, when $12\leq\ell\leq 17$ and $G$ is of type $D_{\ell}$, the second table has to include the weights $\lambda_{1}$ and $\lambda_{2}$ whose associated irreducible representation has dimension $2^{\ell-1}$.
Proof.
The tables in Appendix contain all the information to calculate directly the dimensions of the weights spaces $V(\lambda)$. Thus, we may deduce that the weights in the tables above are effectively admissible. Next we shall prove that the list is exhausted respect to admissible ones. To this end, let
$\lambda=a_{1}\lambda_{1}+a_{2}\lambda_{2}+\cdots+a_{\ell}\lambda_{\ell}$
be a dominant weight and assume that $G$ is of type $C_{\ell}$. By arguing as in proof of Theorem 5.9 (nullity of coefficients) we obtain that $\lambda$ admissible implies
$\lambda=\lambda_{1}\quad\quad\text{ or }\quad\quad\lambda=a_{\ell-3}\lambda_{\ell-3}+a_{\ell-2}\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$
One here finds a little difference in determining the admissible weights with the case $B_{\ell}$, for $\ell=12,13,14,15$. Explicitly, $\lambda=\lambda_{1}$ is not admissible by noticing first $\lambda_{3}\prec\lambda$ and that
$\dim(V(\lambda)\geq|\rm{Orb}_{\mathcal{W}}(\lambda_{3})|=2^{\ell-3}\ell(\ell-1)(\ell-2)>\ell^{4}.$
Thus
$\lambda=a_{\ell-3}\lambda_{\ell-3}+a_{\ell-2}\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{1}\lambda_{\ell}.$
From this point, the rest of the route of $B_{\ell}$ works perfectly for the current case $C_{\ell}$ by resorting to parallel arguments and Appendix -Table 1.
Now, assume that $G$ is of type $D_{\ell}$. By arguing as in proof of Theorem 5.9 too, we obtain that $\lambda$ admissible implies
$\lambda=\lambda_{1},\quad\quad\lambda=\lambda_{2}\quad\quad\text{or}\quad\quad\lambda=a_{\ell-3}\lambda_{\ell-3}+a_{\ell-2}\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$
Contrary to the previous case, here, the weights $\lambda=\lambda_{1}$, $\lambda=\lambda_{2}$ are admissible for $12\leq\ell\leq 17$ since the dimension of $V(\lambda_{i})$ is $2^{\ell-1}$ for $i=1,2$. On the other hand, the dominant weights of the form $\lambda=a_{\ell-3}\lambda_{\ell-3}+a_{\ell-2}\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$ can be addressed as above by using Appendix -Table 2. ∎
Remark 5.11.
It is worth mentioning that:
  • If $G$ is of type $C_{\ell}$ and $V$ the irreducible $G$-module of dimension $2\ell$. Then, the $i$-symmetric power $Sym^{i}(V)$ is irreducible and the anti-symmetric tensor product $\Lambda^{i}(V)$ is reducible for $i=2,3,4$.
  • In contrast with the case $C_{\ell}$, when $G$ is of type $D_{\ell}$ we may observe that $\operatorname{Sym}^{i}(V)$ is reducible and $\Lambda^{i}(V)$ is irreducible for $i=2,3,4$.
6. Some applications
Let $\rho_{p}:G_{\mathbb{Q}}\rightarrow\operatorname{GO}_{n}(\overline{\mathbb{Q}}_{p})\subseteq\operatorname{GL}_{n}(\overline{\mathbb{Q}}_{p})$ be a $p$-adic orthogonal Galois representation. Examples of this kind of Galois representations can be obtained from the so-called regular algebraic, polarizable, cuspidal automorphic representations of $\operatorname{GL}_{n}(\mathbb{A}_{\mathbb{Q}})$, via the global Langlands correspondence [2, §2.1].
In this case, the large image conjecture predicts that, for almost all primes $p$, the image of the proyectivizations $\overline{\rho}^{proj}_{p}:G_{\mathbb{Q}}\rightarrow\operatorname{PGO}_{n}(\overline{\mathbb{F}}_{p})$ of the residual representations $\overline{\rho}_{p}:G_{\mathbb{Q}}\rightarrow\operatorname{GO}_{n}(\overline{\mathbb{F}}_{p})$ should contains $\operatorname{P\Omega}_{n}^{\pm}(\mathbb{F}_{p^{s}})$ for some positive integer $s$. As we pointed out in the introduction, this conjecture is known in very few cases. However, if we impose certain conditions on the local components of $\pi$, at least we can bound the possible images of $\overline{\rho}_{p}$. For example, in [14, Proposition 3.4], the second author proved the following criterion concerning to the image of certain families of orthogonal Galois representations.
Proposition 6.1.
Let =$\mathcal{R}=\{\rho_{p}\}_{p}$ be a family of $p$-adic orthogonal Galois representations $\rho_{p}:G_{\mathbb{Q}}\rightarrow\operatorname{GO}_{n}(\overline{\mathbb{Q}}_{p})$ as in [14, Theorem 2]. Then, for almost all primes $p$, the image of $\overline{\rho}_{p}^{proj}$ is contained in an almost simple group with socle a finite simple group of Lie type in characteristic $p$.
As observed in [14, Section 5], if we only had a best knowledge of the classification of irreducible representations of finite simple groups of Lie type, we could improve Proposition 6.1. In particular, the main result of [14] (Theorem 1.1), states that for almost all primes $p$ and $n=4\varpi$ (where $\varpi$ is a prime number such that $17\leq\varpi\leq 73$), the image of $\overline{\rho}_{p}^{proj}$ is equal to one of the following groups:
(6.1) $\operatorname{P\Omega}^{\pm}_{n}(\mathbb{F}_{p^{s}}),\;\operatorname{PSO}^{\pm}_{n}(\mathbb{F}_{p^{s}}),\;\operatorname{PO}_{n}^{\pm}(\mathbb{F}_{p^{s}}),\;\operatorname{PGO}^{\pm}_{n}(\mathbb{F}_{p^{s}})$
[…]
In the same line of thought, using the strategy in Section 5 of [14] combined with Theorems 4.1, 5.9 and 5.10, we may prove the following improvement of Proposition 6.1 for $n=4\varpi$ with $\varpi$ a prime number such that $79\leq\varpi\leq 787$.
Theorem 6.2.
Let =$\mathcal{R}=\{\rho_{p}\}_{p}$ be a compatible system of $n$-dimensional orthogonal Galois representations $\rho_{p}:G_{\mathbb{Q}}\rightarrow\operatorname{GO}_{n}(\overline{\mathbb{Q}}_{p})$ as in [14, Theorem 2]. If $n=4\varpi$, where $\varpi$ is a prime number such that $79\leq\varpi\leq 787$, then for almost all $p$ the image of $\overline{\rho}_{p}^{proj}$ is equal to
(6.2) $\operatorname{P\Omega}^{\pm}_{n}(\mathbb{F}_{p^{s}}),\;\operatorname{PSO}^{\pm}_{n}(\mathbb{F}_{p^{s}}),\;\operatorname{PO}_{n}^{\pm}(\mathbb{F}_{p^{s}}),\;\operatorname{PGO}^{\pm}_{n}(\mathbb{F}_{p^{s}})$
or to a finite group $PG(\mathbb{F}_{p^{s}})$, where $G$ is of type $A_{\ell}$ ($\ell\leq 7$), $B_{\ell}$ , $C_{\ell}$, $D_{\ell}$ ($\ell\leq 5$) or $G_{2}$, for some positive integer $s$.
Proof.
[…]
Assume that $G$ is of a Lie type different from $A_{\ell}$ ($\ell\leq 7$), $B_{\ell}$ , $C_{\ell}$, $D_{\ell}$ ($\ell\leq 5$) and $G_{2}$. In order to determine all irreducible representations in defining characteristic for groups of a fixed Lie type, we first need to compute all factorizations of $n$ into factors greater than 1. Such factorizations are $4\varpi=4\cdot\varpi=2\cdot 2\cdot\varpi=2\cdot 2\varpi$. Appealing to Tables 6.6-6.53 of [11] and the assumptions on $G$, there are no irreducible representations of dimension 2 or 4. Thus, there are no tensor decomposable representations of dimension $4\varpi$. Moreover, according to Theorems 4.1, 5.9, 5.10 and Tables 6.6-6.53 of [11], the only irreducible tensor indecomposable self-dual representations of dimension $4\varpi$ are the natural representations of $C_{2\varpi}$ and $D_{2\varpi}$. However, the natural representation of $C_{2\varpi}$ is symplectic then the only irreducible $4\varpi$-dimensional orthogonal representation occurs when $G$ is of type $D_{2\varpi}$. ∎
Then, we have the following result:
Corollary 6.3.
For almost all prime $p$ and each $n$ of the form $4\varpi$, where $\varpi$ is a prime number such that $79\leq\varpi\leq 787$, we have that at least one of the following groups:
(6.3) $\operatorname{PSL}_{\ell+1}(\mathbb{F}_{p^{s}}),\;\operatorname{PGL}_{\ell+1}(\mathbb{F}_{p^{s}}),\;\operatorname{PSU}_{\ell+1}(\mathbb{F}_{p^{s}}),\;\operatorname{PGU}_{\ell+1}(\mathbb{F}_{p^{s}})\quad 2\leq\ell\leq 7.$
(6.4) $G_{2}(\mathbb{F}_{p^{s}}),\;\operatorname{P\Omega}_{2\ell+1}(\mathbb{F}_{p^{s}}),\;\operatorname{PSO}_{2\ell+1}(\mathbb{F}_{p^{s}}),\;\operatorname{PSp}_{2\ell}(\mathbb{F}_{p^{s}}),\;\operatorname{PGSp}_{2\ell}(\mathbb{F}_{p^{s}}),\quad 2\leq\ell\leq 5.$
(6.5) $\operatorname{P\Omega}^{\pm}_{2\ell}(\mathbb{F}_{p^{s}}),\;\operatorname{PSO}^{\pm}_{2\ell}(\mathbb{F}_{p^{s}}),\;\operatorname{PO}_{2\ell}^{\pm}(\mathbb{F}_{p^{s}}),\;\operatorname{PGO}^{\pm}_{2\ell}(\mathbb{F}_{p^{s}}),\quad 2\leq\ell\leq 5\mbox{ or }\ell=2\varpi.$
occurs as a Galois group over $\mathbb{Q}$ for infinitely many primes $p$ and infinitely many positive integers $s$.
Remark 6.4.
In the previous result, we include the possible occurrence of (6.3), (6.4), and (6.5), for $2\leq\ell\leq 5$, as Galois groups of $\mathbb{Q}$. In the future, computational tools could be useful to compute all the representations of these groups for dimensions $n\leq 3148$. Moreover, it is worth that the result is new for $\varpi>181$ while for $\varpi\leq 181$ the foregoing result follows from [11].
Remark 6.5.
Finally we also emphasize that if we assume the computations on the web page of Frack Lübek [^1] http://www.math.rwth-aachen.de/ Frank.Luebeck/chev/WMSmall/index.html the Theorem 6.2 works for $100002477\leq\varpi<$ 10000 and for $121553037\leq\varpi<$ 12155 if we include groups of type $C_{6}$. In this case, the result is new for $\varpi>283$ and for $\varpi\leq 283$ follows from Lübek’s computations.
7. Apendix
Next we shall present the multiplicities and lengths of the dominants weights $\mu\preceq\lambda$, for each highest weight $\lambda$ listed in the table in Theorem 5.10. We may remark that these tables provide all the information to compute the dimensions of $V(\lambda)$ which implies admissibility of the foregoing weights $\lambda$.
In what follows we consider orthogonal basis $\{e_{1},e_{2},\cdots,e_{\ell}\}$ of $E\cong\mathbb{R}^{\ell}$.
7.1. Type $C_{\ell}$
Start by considering $\{\alpha_{1},\alpha_{2},\cdots,\alpha_{\ell}\}$ the simple root system of type $C_{\ell}$ given by
$\alpha_{1}:=2e_{\ell}\quad\text{and}\quad\alpha_{i}:=e_{\ell-i+1}-e_{\ell-i+2},\quad\quad 2\leq i\leq\ell,$
We then note that $(\alpha_{1},\alpha_{1})=4$, $(\alpha_{1},\alpha_{2})=-2$ and
(7.1) $(\alpha_{i},\alpha_{j})=\begin{cases}\phantom{-}2,&\text{ if }i=j\\ -1,&\text{ if }i=j\pm 1\\ \phantom{-}0,&\text{ otherwise },\end{cases}$
for $i,j>1$. The fundamental weights $\lambda_{i}$ ($1\leq i\leq\ell$), such that $(\lambda_{i},\alpha_{j})=\delta_{ij}$, are given by
$\lambda_{1}$ $=\alpha_{\ell}+2\alpha_{\ell-1}+\cdots+(\ell-1)\alpha_{2}+\ell(1/2)\alpha_{1}$
$\lambda_{i}$ $=\alpha_{\ell}+2\alpha_{\ell-1}+\cdots+(\ell-i)\alpha_{i+1}+(\ell-i+1)(\alpha_{i}+\cdots+(1/2)\alpha_{1}),\quad 2<i<\ell,$
$\lambda_{\ell}$ $=\alpha_{\ell}+\alpha_{\ell-1}+\cdots\alpha_{2}+(1/2)\alpha_{1}.$
We note that $(\lambda_{i},\lambda_{i})=\ell-i+1$. The half sum of the positive roots is
$\delta=\sum_{i=1}^{\ell}\lambda_{i}=\ell\left(\frac{\ell+1}{2}\right)\cdot\frac{1}{2}\alpha_{1}+\sum_{i=2}^{\ell}(\ell-i+1)\left(\frac{\ell+i}{2}\right)\alpha_{i}.$
It is easy to verify that
$(\alpha_{i},\delta)=\begin{cases}2,&\text{ if }i=1\\ 1,&\text{ otherwise }.\\ \end{cases}$
With the notation above in mind, the table below may be deduced by routing each step of the case $B_{\ell}$.
highest weight dominant weight $\mu\preceq\lambda$ multiplicity of $\mu$ length of $\mathcal{W}$-orbit
$3\lambda_{\ell}$ $3\lambda_{\ell}$ $1$ $2\ell$
$\lambda_{\ell-1}+\lambda_{\ell}$ $1$ $4\ell(\ell-1)$
$\lambda_{\ell-2}$ $1$ $32^{2}\ell(\ell-1)(\ell-2)/$ 3
[…]
$4\lambda_{\ell}$ $4\lambda_{\ell}$ $1$ $2\ell$
$\lambda_{\ell-1}+2\lambda_{\ell}$ $1$ $4\ell(\ell-1)$
$2\lambda_{\ell}$ $1$ $2\ell$
$\lambda_{\ell-2}+\lambda_{\ell}$ $1$ $2^{2}\ell(\ell-1)(\ell-2)$
$\lambda_{\ell-3}$ $1$ $242^{4}\ell(\ell-1)(\ell-2)(\ell-3)/$ 24
[…]
00 $\ell(\ell+1)/2$ $1$
$\lambda_{\ell-2}$ $\lambda_{\ell-2}$ $1$ $62^{3}\ell(\ell-1)(\ell-2)/$ 6
$\lambda_{\ell}$ $\ell-2$ $2\ell$
$\lambda_{\ell-3}$ $\lambda_{\ell-3}$ $1$ $242^{4}\ell(\ell-1)(\ell-2)(\ell-3)/$ 24
$\lambda_{\ell-1}$ $\ell-3$ $62\ell(\ell-1)/$ 6
00 $\ell(\ell-3)/2$ $1$
$\lambda_{\ell-1}+\lambda_{\ell}$ $\lambda_{\ell-1}+\lambda_{\ell}$ $1$ $2^{2}\ell(\ell-1)$
$\lambda_{\ell-2}$ $2$ $32^{2}\ell(\ell-1)(\ell-2)/$ 3
$\lambda_{\ell}$ $2(\ell-1)$ $2\ell$
Table 1.
7.2. Type $D_{\ell}$
Start by considering $\{\alpha_{1},\alpha_{2},\cdots,\alpha_{\ell}\}$ the simple root system of type $D_{\ell}$ given by
$\alpha_{1}:=e_{\ell-1}+e_{\ell}\quad\text{and}\quad\alpha_{i}:=e_{\ell-i+1}-e_{\ell-i+2},\quad\quad 2\leq i\leq\ell.$
We then note that $(\alpha_{1},\alpha_{1})=2$, $(\alpha_{1},\alpha_{2})=0$, $(\alpha_{1},\alpha_{3})=-1$ and
(7.2) $(\alpha_{i},\alpha_{j})=\begin{cases}\phantom{-}2,&\text{ if }i=j\\ -1,&\text{ if }i=j\pm 1\\ \phantom{-}0,&\text{ otherwise },\\ \end{cases}$
for $i,j>1$. The fundamental weights $\lambda_{i}$ ($1\leq i\leq\ell$), such that $(\lambda_{i},\alpha_{j})=\delta_{ij}$, are given by
$\lambda_{1}$ $=(1/2)(\alpha_{\ell}+2\alpha_{\ell-1}+\cdots+(\ell-2)\alpha_{3}+(1/2)(\ell-2)\alpha_{2}+(1/2)\ell\alpha_{1})$
$\lambda_{2}$ $=(1/2)(\alpha_{\ell}+2\alpha_{\ell-1}+\cdots+(\ell-2)\alpha_{3}+(1/2)\ell\alpha_{2}+(1/2)(\ell-2)\alpha_{1}),$
$\lambda_{i}$ $=\alpha_{\ell}+2\alpha_{\ell-1}+\cdots+(\ell-i)\alpha_{i+1}+(\ell-i+1)(\alpha_{i}+\alpha_{i-1}+\cdots+\alpha_{3})+(1/2)(\ell-i+1)(\alpha_{2}+\alpha_{1}),\quad 2<i<\ell,$
$\lambda_{\ell}$ $=\alpha_{\ell}+\alpha_{\ell-1}+\cdots+\alpha_{3}+(1/2)(\alpha_{2}+\alpha_{1}).$
We observe that $(\lambda_{i},\lambda_{i})=\ell-i+1$ for $3\leq i\leq\ell$ and $(\lambda_{1},\lambda_{1})=(\lambda_{2},\lambda_{2})=\ell/4$. The half sum of the positive roots is
$\delta$ $=\frac{1}{4}\ell(\ell-1)(\alpha_{1}+\alpha_{2})+\frac{1}{2}\sum_{i=3}^{\ell}(\ell-i+1)(\ell+i-2)\alpha_{i}.$
It is easy to verify that
$(\alpha_{i},\delta)=\begin{cases}2&\text{ if }i=1,2\\ 1&\text{ otherwise }.\\ \end{cases}$
With the notation above in mind, the table below may be deduced as in the previous subsection.
highest weight dominant weight $\mu\preceq\lambda$ multiplicity of $\mu$ length of $\mathcal{W}$-orbit
$3\lambda_{\ell}$ $3\lambda_{\ell}$ $1$ $2\ell$
$\lambda_{\ell-1}+\lambda_{\ell}$ 1 $4\ell(\ell-1)$
$\lambda_{\ell-2}$ 1 $32^{2}\ell(\ell-1)(\ell-2)/$ 3
$\lambda_{\ell}$ $\ell-1$ $2\ell$
$4\lambda_{\ell}$ $4\lambda_{\ell}$ $1$ $2\ell$
$\lambda_{\ell-1}+2\lambda_{\ell}$ $1$ $4\ell(\ell-1)$
$2\lambda_{\ell-1}$ $1$ $2\ell(\ell-1)$
$\lambda_{\ell-2}+\lambda_{\ell}$ $1$ $2^{2}\ell(\ell-1)(\ell-2)$
$\lambda_{\ell-3}$ $1$ $242^{2}\ell(\ell-1)(\ell-2)(\ell-3)/$ 24
[…]
00 $\ell(\ell-1)/2$ 1
$\lambda_{\ell-2}$ $\lambda_{\ell-2}$ $1$ $62^{3}\ell(\ell-1)(\ell-2)/$ 6
$\lambda_{\ell}$ $\ell-1$ $2\ell$
$\lambda_{\ell-3}$ $\lambda_{\ell-3}$ $1$ $242^{4}\ell(\ell-1)(\ell-2)(\ell-3)/$ 24
$\lambda_{\ell-1}$ $\ell-2$ $2\ell(\ell-1)$
00 $\ell(\ell-1)/2$ 1
$\lambda_{\ell-1}+\lambda_{\ell}$ $\lambda_{\ell-1}+\lambda_{\ell}$ $1$ $2^{2}\ell(\ell-1)$
$\lambda_{\ell-2}$ $2$ $62^{3}\ell(\ell-1)(\ell-2)/$ 6
$\lambda_{\ell}$ $2(\ell-1)$ $2\ell$
Table 2.